### v1

# Regularization methods for solving hierarchical variational inequalities with complexity guarantees

Abstract
We consider hierarchical variational inequality problems, or more generally, variational inequalities defined over the set of zeros of a monotone operator. This framework includes convex optimization over equilibrium constraints and equilibrium selection problems. In a real Hilbert space setting, we combine a Tikhonov regularization and a proximal penalization to develop a flexible double-loop method for which we prove asymptotic convergence and provide rate statements in terms of gap functions. Our method is flexible, and effectively accommodates a large class of structured operator splitting formulations for which fixed-point encodings are available. Finally, we validate our findings numerically on various examples.
1 Introduction
Bi-level optimization problems consist of two nested optimization formulations, referred to as the inner and the outer problem. Solutions to the inner problem determine the feasible set over which a minimizer of an outer objective function is searched for. This class of problems is an extremely active area of research recently, driven by important practical applications in engineering, economics and machine learning. For recent comprehensive reviews, see [dempe2020bilevel, liu2021investigating]. A particularly appealing class of bi-level optimization problems are those in which only a single decision variable is involved at both layers of the problem. These special cases are often called simple bi-level optimization problems [Dempe:2021aa]. In this special instance, the problem is given by the convex constrained minimization problem
$\min_{x\in\mathcal{S}_{0}}g(x),\qquad\text{where}~\mathcal{S}_{0}=\operatorname*{argmin}_{x\in\mathcal{H}}\{f(x)+r(x)\},$
$\mathcal{S}_{0}$ is the set of minimizers of a the inner level problem, where $f$ and $g$ are convex and continuously differentiable and $r$ is an extended-valued proper closed convex function. This problem received significant attention over the last years, with important advances made in [beck2014first, SabSht17, MerchavSIOPT23, AminiYous19, shen2023online] and [yousefian2021bilevel]. Moving beyond the convex optimization formulation of the inner and the outer level problem, recent papers consider a more general problem formulation by replacing them by variational inequalities. While this is a classical formulation in inverse problems and signal processing [deutsch1998minimizing, marino2006general, Yamada2011MinimizingTM], only a few papers obtain rates of convergence in terms of merit functions. Motivated by narrowing this gap, we consider the problem of solving a class of constrained variational inequality problems
$\operatorname{VI}(\mathsf{G},\mathcal{S}_{0})\colon\text{ Find $u\in\mathcal{S}_{0}$ s.t. }\langle\mathsf{G}u,v-u\rangle\hskip 17.00024pt\forall v\in\mathcal{S}_{0}=\operatorname{Zer}(\mathsf{M}),$ (P)
where $\mathsf{G}\colon\mathcal{H}\to\mathcal{H}$ is a monotone and Lipschitz continuous operator, and $\mathcal{S}_{0}=\operatorname{Zer}(\mathsf{M})$ is the (non-empty) set of zeros of another maximally monotone operator $\mathsf{M}\colon\mathcal{H}\to 2^{\mathcal{H}}$, all defined on a real Hilbert space $\mathcal{H}$ with inner product $\langle\cdot,\cdot\rangle$ and corresponding norm $\|{\cdot}\|$. Besides the simple bi-level optimization problem mentioned above, the family of problems (P) contains many specific instances heavily studied in the literature.
Example 1.1 (Hierarchical Variational Inequality).
Given a mapping $\mathsf{F}\colon\mathcal{H}\to\mathcal{H}$ and a closed convex set $\mathcal{K}\subseteq\mathcal{H}$, the variational inequality problem $\operatorname{VI}(\mathsf{F},\mathcal{K})$, is the problem of finding a point $u\in\mathcal{K}$ such that
$\langle\mathsf{F}(u),v-u\rangle\geq 0\hskip 17.00024pt\forall v\in\mathcal{K}.$
The solution set of $\operatorname{VI}(\mathsf{F},\mathcal{K})$ can be expressed as $\operatorname{Zer}(\mathsf{F}+\operatorname{\mathsf{NC}}_{\mathcal{K}})$, involving the normal cone operator
$\operatorname{\mathsf{NC}}_{\mathcal{K}}(v)\triangleq\left\{\begin{array}[]{ll}\varnothing&\text{if }v\notin\mathcal{K}\\ \{p\in\mathcal{H}|\langle p,u-v\rangle\leq 0,\;\forall u\in\mathcal{K}\}&\text{ else.}\end{array}\right$.
Hence, if we set the operator $\mathsf{M}=\mathsf{F}+\operatorname{\mathsf{NC}}_{\mathcal{K}}$, Problem (P) specializes to a hierarchical variational inequality problem given by
$\operatorname{VI}(\mathsf{G},\mathcal{S}_{0}),\qquad\text{where $\mathcal{S}_{0}$ is the solution set of $\operatorname{VI}(\mathsf{F},\mathcal{K})$}.$
For recent studies of such problems in finite-dimensional euclidean spaces, we refer to [alves_inertial_2025] and [samadi_improved_2025].
Example 1.2 (Equilibrium Selection).
A mathematical program with equilibrium constraints [Luo_Pang_Ralph_1996] optimizes a real-valued function over a feasible set described as the solution set of an equilibrium problem. As such, one can select amongst lower-level equilibria via a convex design criterion $g\colon\mathcal{H}\to\mathbb{R}$ by writing
$\min_{u\in\mathcal{H}}\left\{g(u)~~\colon~~u\in\mathcal{S}_{0}\right\}.$
If $g$ is convex and Fréchet differentiable, this becomes a hierarchical variational inequality with $\mathsf{G}=\nabla g$, hence a particular instance of Problem (P). This viewpoint covers a broad range of models encountered in, e.g., engineering, economics, inverse problems and control (see [hintermuller2011first, hintermuller2015bi-level, hintermuller2014several, outrata2000generalized, de2023bi-level]).
1.1 Contributions
Numerical schemes with complexity statements for resolving Problem (P) have been recently published in the papers [Kaushik:2021aa, samadi_improved_2025] and [alves_inertial_2025]. [Kaushik:2021aa] extend the simple bi-level setting to a lower level problem defined in terms of a variational inequality while keeping the upper level problem as a convex optimization problem. This is a special instance of Problem (P), which is related to mathematical programming with equilibrium constraints [Luo_Pang_Ralph_1996], and to the equilibrium selection problem in game theory [Benenati:2022aa, Benenati:2023aa]. The recent papers [samadi_improved_2025, alves_inertial_2025] go one step further and study the special instance of Problem (P) in which the upper and lower level problem are defined in terms of a variational inequality problem.
In this paper we generalize the aforementioned existing results by considering a more general class of hierarchical equilibrium problems admitting the formulation (P). We prove rates on both levels of the equilibrium problem via a very flexible numerical scheme, which can easily be modulated with respect to problem structure (specifically of the lower-level inclusion problem). Our main contributions are threefold:
  • Hilbert space setting. We extend analyses that are typically confined to finite-dimensional Euclidean spaces to possibly infinite-dimensional real Hilbert spaces.
  • Monotone-inclusion lower level and modular oracle. Rather than considering a variational inequality at the lower level, we allow the feasible set to be defined as the solution set of an abstract monotone inclusion. This formulation allows to draw from the rich theory of monotone operators and enables a very general algorithmic design based on a monotone-inclusion oracle obtained through a fixed-point encoding, which can be tailored to the structure of the operators involved.
  • Quantitative guarantees via gap functions. While much of the existing literature focuses on weak convergence results without explicit rate statements, we introduce gap functions for both the lower- and upper-level problems and establish quantitative convergence guarantees for these measures. Our results are formulated within a specific geometric setting in which we assume that the feasibility gap function is assumed to satisfy an Hölderian error bound condition such as [Cabot2005] in the context of hierarchical optimization and more recently [samadi_improved_2025] and [boct2025accelerating].
We achieve these results via the careful design of an algorithmic template drawing inspiration from classical diagonal methods involving iterative Tikhonov and proximal regularization via anchor terms. By constructing a family of auxiliary combined equilibrium problems, our method constructs a sequence in an inner loop method, which is designed to track a sequence of temporal solutions. This procedure is essentially an inertial Krasnoselskii-Mann iteration and thus only relies on the availability of a ready-to-compute fixed-point encoding map for the lower-level solution set $\mathcal{S}_{0}=\operatorname{Zer}(\mathsf{M})$. An outer-loop procedure updates the Tikhonov parameter and the anchor term to steer the trajectory of the numerical algorithm towards a solution of the entire Problem (P).
Diagonal schemes for solving convex programs are a classical topic [alart1991penalization]. The idea of extending this approach to hierarchical equilibrium problems is taken from [Facchinei:2014aa] and [Lampariello:2020aa]. In particular, [Lampariello:2020aa] also proposes a kind of double-loop scheme with a forward-backward iteration acting in the inner loop. In this paper, we consider a more general class of splitting problems, with explicit rate guarantees at both levels of the problem.
Specifically, in the case where $\mathsf{G}$ is merely monotone (opposed to strongly monotone), we obtain weak convergence of the averaged sequence of iterates $(\overline{w}_{N})$, with rates of the form, for arbitrary $b\in(0,1)$,
$\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{N})\leq\tilde{\mathcal{O}}\left(\frac{1}{N^{1-b}}\right),\qquad\text{and}\qquad 0\leq\operatorname{Gap}_{\operatorname{feas}}(\overline{w}_{N})\leq\tilde{\mathcal{O}}\left(\frac{1}{N^{\min(b,1-b)}}\right),$
for suitably defined optimality and feasibility gaps (see Section 3.2 for precise definition). In this setting, $\tilde{\mathcal{O}}$ disregards logarithmic factors, which are natural consequences of the double-loop structure of our algorithm. Ignoring these logarithmic factors, which previous works do not obtain due to single-loop structures, our work is on par with established benchmarks:
  • In [Kaushik:2021aa, Corollary 3.5], the authors study optimization problems with Cartesian variational inequality constraints, where the objective is convex and the constraint set is a Cartesian product associated with a monotone mapping. Their algorithm resembles a forward-backward method, cast into a single-loop scheme. For the deterministic variant of their method, they obtain rates of $\mathcal{O}(1/N^{0.5-b})$ for the optimality gap and $\mathcal{O}(1/N^{b})$ for the feasibility gap, for $b\in(0,0.5)$. Therefore, ignoring the logarithmic factors, the rates achieved in the present work are more favorable.
  • In [samadi_improved_2025, Theorem 3.6], the authors study a regularized extragradient method, and establish rates of $\mathcal{O}(1/N^{1-b})$ for the optimality gap and $\mathcal{O}(1/N^{b})$ for the feasibility gap. In the $b\in(0,0.5]$ regime our rates match with theirs, and we improve upon their rates in the $b\in(0.5,1)$ regime. We do again recall that our rates include logarithmic terms that their avoids. We note that [alves_inertial_2025, Theorem 3.8] studies an inertial version of the same method, but does not improve upon the rates.
  • The recent work [marschner_tikhonov_2025, Theorem 9] analyzes a special case of our general algorithmic template, focusing on the forward-backward splitting only. The analysis provided is weaker, in particular, no rate statements are contained in that work.
When $\mathsf{G}$ is strongly monotone, our rates improve to
$\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{N})\leq\tilde{\mathcal{O}}\left(\frac{1}{N}\right),\qquad\text{and}\qquad 0\leq\operatorname{Gap}_{\operatorname{feas}}(\overline{w}_{N})\leq\tilde{\mathcal{O}}\left(\frac{1}{N}\right).$
These rates are comparable to those in [samadi_improved_2025, Theorem 4.6], where a regularized extragradient method is analyzed. To the best of our knowledge, no other works address hierarchical programming problems in the strongly monotone setting in a manner that allows a direct comparison.
Organization of the paper.
The remainder of the paper is organized as follows. The rest of this section includes notation and definitions. A general presentation of the double loop scheme is presented in Section 2, and its analysis is given in Section 3. Section 4 gives examples for the fixed-point encoding map used in the inner loop of our method. Numerical experiments are then performed in Section 5. We conclude in Section 6.
1.2 Notation and Definitions
Let $\mathcal{H}$ be a real Hilbert space with inner product $\langle\cdot,\cdot\rangle$ and corresponding norm $\|{\cdot}\|$. A set-valued operator $\mathsf{A}\colon\mathcal{H}\to 2^{\mathcal{H}}$ is monotone if $\langle u-v,b-a\rangle\geq 0$ for all $(u,b),(v,a)\in\operatorname{Gr}(\mathsf{A})\triangleq\{(v,a)\in\mathcal{H}\times\mathcal{H}|a\in\mathsf{A}(v)\}$. The operator $\mathsf{A}\colon\mathcal{H}\to 2^{\mathcal{H}}$ is maximally monotone if it is monotone and there exists no other monotone operator whose graph properly contains $\operatorname{Gr}(\mathsf{A})$. The resolvent of an operator $\mathsf{A}$ is defined as $\operatorname{\mathsf{J}}_{\mathsf{A}}=(\operatorname{Id}+\mathsf{A})^{-1}$. If $\mathsf{A}$ is maximally monotone, then the resolvent is a nonexpansive operator and thus single-valued. A mapping $\mathsf{T}\colon\mathcal{H}\to\mathcal{H}$ is quasi-contractive if $\|Tx-p\|\leq q\|x-p\|$ for all $p\in\text{Fix}(T)$ and some $q\in(0,1)$. We refer to [BauCom16] for more details on operator theory.
For two sequences $(a_{n})\subset\mathbb{R}$ and $(b_{n})\subset\mathbb{R}$, we say that $a_{n}=O(b_{n})$ if there exists a constant $C>0$ such that $a_{n}\leq Cb_{n}$ for $n$ sufficiently large. We say that $a_{n}=o(b_{n})$ if $\lim_{n\to\infty}{b_{n}}/{a_{n}}=0$. Finally, we say that $a_{n}=\Theta(b_{n})$ if there exists a constant $C>0$ such that $\lim_{n\to\infty}{b_{n}}/{a_{n}}=C$.
2 Outline of the Algorithm
In this section, we first formulate the main assumptions imposed on Problem (P).
Assumption 1.
We assume the following on Problem (P):
  1. 1.
    $\mathsf{G}:\mathcal{H}\to\mathcal{H}$ is a $\mu$-strongly monotone, with $\mu\geq 0$, and $L_{\mathsf{G}}$-Lipschitz continuous operator;
  2. 2.
    $\mathsf{M}\colon\mathcal{H}\to 2^{\mathcal{H}}$ is maximally monotone with bounded domain;
  3. 3.
    The set $\mathcal{S}_{0}$ is nonempty.
Remark 2.1.
Assumption 1 guarantees that the overall Problem (P) has a solution. In particular, $\mathcal{S}_{0}$ is a convex and compact subset of $\mathcal{H}$ [BauCom16].
Remark 2.2.
We emphasize that our assumption includes the case $\mu=0$, in which $\mathsf{G}$ is only Lipschitz and monotone.
For given parameters $(\alpha,\beta)\in(0,\infty)^{2}$ and an anchor point $w\in\mathcal{H}$, we define the associated auxiliary problem as finding
$\overline{u}_{\beta}(w)\in\operatorname{Zer}(\mathsf{M}+\beta\mathsf{G}+\alpha(\operatorname{Id}-w)).$ (Aux)
Here, $\alpha>0$ is a proximal parameter enforcing strong monotonicity, ensuring that the point $\overline{u}_{\beta}(w)$ is the unique zero of $\mathsf{M}+\beta\mathsf{G}+\alpha(\operatorname{Id}-w)$. The parameter $\beta>0$ is a Tikhonov parameter that balances the importance of the lower- and upper-level problems.
Our numerical scheme consists of two main procedures. The inner procedure is an inertial Krasnoselskii-Mann (KM) iteration, whose last iterate is designed to lie in a sufficiently small ball around the temporal solution $\overline{u}_{\beta}(w)$, whose radius is part of the parameters of the procedure. We use this last iterate as a new anchor point, and restart the KM iteration with fresh parameters. To this end, we assume access to a family of parameterized fixed-point mappings $\mathsf{T}_{k}\equiv\mathsf{T}^{(w,\beta)}_{k}\colon\mathcal{H}\to\mathcal{H}$, which satisfy the following condition.
Assumption 2.
For all $k\in\mathbb{N}$ and $(w,\beta)\in\mathcal{H}\times\mathbb{R}_{++}$, the operator $\mathsf{T}^{(w,\beta)}_{k}$ is a quasi-contraction with parameter $q_{k}\in(0,\overline{q})$ for some $\overline{q}\in(0,1)$, sharing a common fixed point $p(w,\beta)$ for all $k\in\mathbb{N}$. Moreover, $\mathsf{T}_{k}^{(w,\beta)}$ is related to Problem (Aux) via a nonexpansive fixed-point transportation map $\mathsf{Z}^{(w,\beta)}\colon\mathcal{H}\to\mathcal{H}$ such that $\overline{u}_{\beta}(w)=\mathsf{Z}^{(w,\beta)}(p(w,\beta))$.
Concrete examples of such operators are provided in Section 4. Generally, the operators may vary over the inner loop as a result of a varying step-size in the selected method.
Remark 2.3.
The proximal regularization parameter $\alpha>0$ is treated in our approach as an external input to the algorithm. Hence, its concrete numerical value used in the computation affects the central funnel $\overline{u}_{\beta}(w)$, as well as the fixed-point encoding map $\mathsf{T}^{(w,\beta)}_{k}$ and eventually also the fixed-point transportation map $\mathsf{Z}^{(w,\beta)}$. However, since it is a fixed parameter, we simplify the notation and omit an explicit dependence from the involved mappings and operators on the value of $\alpha$.
2.1 Inner Loop
The inner loop of our tracking algorithm employs the procedure $\mathtt{KM}(v,\mathsf{T}^{(w,\beta)}_{\bullet},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$, see Algorithm 1, which is an inertial Krasnoselskii-Mann iteration with user-provided parameters embodied in terms of the initial point $v\in\mathcal{H}$, fixed-point encoding map $\mathsf{T}^{(w,\beta)}_{\bullet}$ satisfying Assumption 2, as well as momentum parameters $(\tau_{k})_{k}$ and relaxation parameters $(\theta_{k})_{k}$. This fixed-point iteration is designed to return an approximation of the central funnel fixed-point $\overline{u}_{\beta}(w)$ via the fixed-point transportation map $\mathsf{Z}^{(w,\beta)}$.
Algorithm 1 Function $\mathtt{KM}(v,\mathsf{T}^{(w,\beta)}_{\bullet},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$
$v_{1}=v_{0}:=v$
Set $k=1$
while $\|v_{k+1}-z_{k}\|>\varepsilon$ do
   $k\leftarrow k+1$
   $z_{k}\leftarrow v_{k}+\tau_{k}(v_{k}-v_{k-1})$
   $v_{k+1}\leftarrow(1-\theta_{k})z_{k}+\theta_{k}\mathsf{T}^{(w,\beta)}_{k}(z_{k})$
end while
return $v_{k+1}$
Algorithm 1 employs the stopping time
$\mathbb{K}(\varepsilon)\triangleq\inf\{k\geq 1|\;\|{v_{k+1}-z_{k}}\|\leq\varepsilon\}.$ (Stop)
This stopping criterion is motivated by the following approximation result.
Lemma 2.1.
Let $(\tau_{k})_{k}\subset[0,1],(\theta_{k})_{k}\subset[0,\overline{\theta}]$ for $\overline{\theta}\in(0,1)$, $\varepsilon>0$, and let $(\mathsf{T}_{k})_{k}$ be a sequence of operators satisfying Assumption 2. Denote by $p$ the common fixed-point of $(\mathsf{T}_{k})_{k}$. If $\mathbb{K}(\varepsilon)$ is finite and $v_{\mathbb{K}(\varepsilon)+1}$ is the result of Algorithm 1, then
$\|{v_{\mathbb{K}(\varepsilon)+1}-p}\|\leq\mathtt{e}\triangleq\frac{\varepsilon}{\overline{\theta}}\left({1-\overline{\theta}}+\frac{\overline{q}}{1-\overline{q}}\right).$ (2.1)
Proof.
Given $\varepsilon>0$, denote $\mathbb{K}=\mathbb{K}(\varepsilon)$. By definition of $v_{k+1}$ and quasi-contractiveness of $\mathsf{T}_{k}$, we get
$\|{z_{k}-p}\|$ $=\|{z_{k}-\mathsf{T}_{k}(z_{k})+\mathsf{T}_{k}(z_{k})-p}\|\leq\|{\frac{v_{k+1}-z_{k}}{\theta_{k}}}\|+\|{\mathsf{T}_{k}(z_{k})-p}\|$
$\leq\frac{\varepsilon}{\theta_{k}}+\overline{q}\cdot\|{z_{k}-p}\|.$
Whence $\|{z_{k}-p}\|\leq\frac{\varepsilon}{\theta_{k}(1-\overline{q})}$. We conclude that
$\|{v_{\mathbb{K}+1}-p}\|$ $=\|{v_{\mathbb{K}+1}-\mathsf{T}_{k}(z_{\mathbb{K}})+\mathsf{T}_{k}(z_{\mathbb{K}})-p}\|$
$\leq\|{\frac{(1-\theta_{\mathbb{K}})(z_{\mathbb{K}}-v_{\mathbb{K}+1})}{\theta_{\mathbb{K}}}}\|+\|{\mathsf{T}_{k}(z_{\mathbb{K}})-p}\|$
$\leq\varepsilon\frac{1-\theta_{\mathbb{K}}}{\theta_{\mathbb{K}}}+\overline{q}\cdot\|{z_{\mathbb{K}}-p}\|$
$\leq\frac{\varepsilon}{\theta_{\mathbb{K}}}\left({1-\theta_{\mathbb{K}}}+\frac{\overline{q}}{1-\overline{q}}\right)\leq\frac{\varepsilon}{\overline{\theta}}\left({1-\overline{\theta}}+\frac{\overline{q}}{1-\overline{q}}\right),$
as wanted. $\blacksquare$
Lemma 2.1 has important consequences for the tracking properties of the anchor points $w$ which we recursively generate with the restarting procedure explained next. To give an outlook, consider the sequence $(v_{k})_{k}$ produced by the function $\mathtt{KM}(v,\mathsf{T}^{(w,\beta)}_{\bullet},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$, and let $w^{+}\triangleq\mathsf{Z}^{(w,\beta)}(v_{\mathbb{K}(\varepsilon)+1})$. Lemma 2.1 guarantees that this point satisfies the tracking guarantee
$\|{w^{+}-\overline{u}_{\beta}(w)}\|\leq\mathtt{e}=\frac{\varepsilon}{\overline{\theta}}\left({1-\overline{\theta}}+\frac{\overline{q}}{1-\overline{q}}\right),$ (2.2)
by nonexpansiveness of $\mathsf{Z}^{(w,\beta)}$. The upper bound is thus of order $O(\varepsilon)$, and thus controlled by the user-defined targeted precision.
2.2 Restarting Procedure
Given an anchor point $w$, Algorithm 1 produces a sequence of points $(v_{k})$ whose last iterate corresponds to the output of the function $\mathtt{KM}(v,\mathsf{T}^{(w,\beta)}_{\bullet},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$ and which provides a good candidate for a new anchor point, given the tracking property (2.2). Repeating this over time for a fixed number of iterations is the key updating step constituting our outer loop procedure. The mechanics of this procedure gives rise to an diagonal equilibrium tracking method (DANTE), formally described in Algorithm 2.
Algorithm 2 DANTE (DiAgoNal equilibrium Tracking mEthod)
$w_{0}\in\mathcal{H}$ and $(\tau_{0,k}),(\theta_{0,k})_{k}$ given.
Set $\lambda_{0}=1$ and $S_{0}=0$.
Set $\overline{w}_{0}=w_{0}$.
for $n=0,\ldots,N-1$ do
   Set $v=\mathtt{KM}(w_{n},\mathsf{T}^{(w_{n},\beta_{n})}_{n,\bullet},(\tau_{n,k})_{k},(\theta_{n,k})_{k},\varepsilon_{n})$.
   Let $w_{n+1}=\mathsf{Z}^{(w_{n},\beta_{n})}(v)$ denote the next anchor point.
   Update $\lambda_{n+1}=\lambda_{n}(1+\frac{2\mu\beta_{n}}{\alpha})$ and $S_{n+1}=S_{n}+\lambda_{n}\beta_{n}$
   Update
$\overline{w}_{n+1}=\frac{S_{n}\overline{w}_{n}+\lambda_{n}\beta_{n}w_{n+1}}{S_{n+1}}.$
   Compute $\mathtt{e}_{n}$ based on Equation (2.1).
   Update $\beta_{n+1},(\tau_{n+1,k})_{k},(\theta_{n+1,k})_{k},\varepsilon_{n+1}$.
end for
return $\overline{w}_{N}$
Note that DANTE is a method rather than an algorithm as we do not specify how the parameters required by the fixed-point iteration procedure $\mathtt{KM}(v,\mathsf{T}^{(w,\beta)}_{\bullet},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$ are set. This is a part of the user-defined input and should be decided based on the eventually known problem structure.
3 Convergence Analysis
In this section, we focus on the convergence analysis of our proposed method. The double-loop structure naturally splits the analysis into the analysis of the inner loop (Section 3.1), the analysis of the restarting procedure (Section 3.2), and the analysis of how they interact (Section 3.3).
3.1 Convergence Analysis of Inner Loop
Before proving that Algorithm 1 terminates in finite time, we recall the following assumption from [Cortild:2024aa]. Note that the conditions for Lemma 2.1 are verified under Assumptions 2 and 3, when considering a single outer iteration.
Assumption 3.
[Cortild:2024aa, Hypothesis 2.1] Assume $(\theta_{k})_{k}\subset(0,\overline{\theta}]$ and that $(\tau_{k})_{k}\subset[0,\overline{\tau}]$ is monotonically non-decreasing, where $\overline{\theta},\overline{\tau}\in(0,1)$. Moreover, assume that, for all $k\in\mathbb{N}$,
$Q_{k}\tau_{k}(1+\tau_{k})+(\theta_{k}^{-1}-1)\tau_{k}(1-\tau_{k})-Q_{k}(\theta_{k-1}^{-1}-1)(1-\tau_{k-1})<0,$ (3.1)
where $Q_{k}=1-\theta_{k}(1-q_{k}^{2})$, and that $\overline{Q}\triangleq\sup_{k\in\mathbb{N}}Q_{k}<1$.
Under the above assumption, we can guarantee strong convergence of the iterates.
Theorem 3.1.
[Cortild:2024aa, Theorem 2.2] Assume Assumptions 2 and 3 hold. If $(z_{k},v_{k})_{k}$ is generated by Algorithm 1 without stopping criterion, then $(z_{k},v_{k})_{k}$ converges strongly to $(p,p)=(\textnormal{Fix}(\mathsf{T}_{k}),\textnormal{Fix}(\mathsf{T}_{k}))$. Moreover $\sum_{k\geq 1}\|{v_{k}-p}\|^{2}<\infty$, and specifically
$\|{v_{k}-p}\|^{2}\leq\frac{\overline{Q}^{k}}{(1-\overline{\tau})(1-\overline{\theta})}\cdot\|{v_{0}-p}\|^{2}$
for all $k\geq 1$.
An immediate corollary is that the stopping criterion in Algorithm 1 allows to conclude finite-time termination.
Corollary 3.2.
Assume Assumptions 2 and 3 hold. It holds that $\sum_{k\geq 1}\|v_{k+1}-z_{k}\|^{2}<\infty$. Specifically, $\|v_{k+1}-z_{k}\|\to 0$, and hence Algorithm 1 terminates in finite time. Moreover, it holds that
$\|{v_{k+1}-z_{k}}\|^{2}\leq\overline{Q}^{k-1}\cdot\frac{4(1+\overline{\tau})^{2}}{(1-\overline{\tau})(1-\overline{\theta})}\cdot\|{v_{0}-p}\|^{2}$
for all $k\geq 2$.
Proof.
Note that
$\|v_{k+1}-z_{k}\|$ $\leq\|v_{k+1}-p\|+\|z_{k}-p\|$
$\leq\|v_{k+1}-p\|+(1+\tau_{k})\|v_{k}-p\|+\tau_{k}\|v_{k-1}-p\|$
$\leq 2(1+\overline{\tau})\max(\|v_{k+1}-p\|,\|v_{k}-p\|,\|v^{k-1}-p\|).$
The result follows by the square-summability of $(v_{k}-p)$, and the rate provided in Theorem 3.1. $\blacksquare$
The rate in Corollary 3.2 does not provide a user-friendly quantitative result unless one can control the quantity $\|v_{0}-p\|$. We present an additional set of assumptions under which this can be done.
Assumption 4.
We assume that $\operatorname{dom}(\mathsf{T}_{k})\subset\operatorname{dom}(\mathsf{M})$ for all $k\in\mathbb{N}$.
Remark 3.1.
Assumption 4 might seem restrictive, but in fact, they are satisfied by many operators. For instance, if $\mathsf{M}=\mathsf{F}+\operatorname{\mathsf{NC}}_{\mathcal{K}}$, where $\operatorname{\mathsf{NC}}_{\mathcal{K}}$ is the normal cone of ${\mathcal{K}}=\operatorname{dom}(\mathsf{M})$ and $\mathsf{F}$ is a Lipschitz monotone operator, the forward-backward operator (FB) satisfies the desired conditions, as the resolvent of $\operatorname{\mathsf{NC}}_{\mathcal{K}}$ reduces to a projection onto $\mathcal{K}$}, such that $\mathsf{T}(x)\in{\mathcal{K}}$ for all $x\in\mathcal{H}$. The forward-backward operator and other examples are discussed in Section 4.
Under these assumptions, we can achieve a user-friendly rate, as presented in the following corollary.
Corollary $3.$
Assume Assumptions 2, 3, and 4 hold. Then
$\|v_{k+1}-z_{k}\|\leq C\cdot\overline{Q}^{k/2},$
for all $k\geq 1$, where
$C\triangleq\frac{2(1+\overline{\tau})}{\sqrt{\overline{Q}(1-\overline{\tau})(1-\overline{\theta})}}\cdot\operatorname{diam}(\operatorname{dom}(\mathsf{M})).$ (Def-$C$)
Specifically,
$\mathbb{K}(\varepsilon)\leq\overline{K}_{\varepsilon}\triangleq\left\lceil 2\log(C/\varepsilon)/\log(1/\overline{Q})\right\rceil.$
Proof.
The conditions guarantee that all iterates are in $\operatorname{dom}(\mathsf{M})$, thus justifying the bounds. Moreover, since $\overline{Q}\in(0,1)$, $C\cdot\overline{Q}^{k/2}\leq\varepsilon$ implies the bound on $k$. $\blacksquare$
Remark 3.2.
By [Cortild:2024aa, Theorem 2.1], we also know that weak convergence of the iterates $(v_{k})$ and the velocities $z_{k}$ is guaranteed when $\mathsf{T}_{k}$ is just nonexpansive. Moreover, if $\mathsf{T}_{k}$ is a fixed-point encoding of Problem (Aux), which has a unique solution $p$ by construction, we know that the sequence convergences weakly to this unique fixed-point $p$. Specifically, in the finite-dimensional setting, we know that $\|v_{k}-p\|\to 0$ and $\|v_{k+1}-z_{k}\|\to 0$, meaning that Algorithm 2 still can be applied. It is however important to note that Lemma 2.1 no longer holds, and that we do not have an appropriate stopping criterion to guarantee $\|w^{+}-\overline{u}_{\beta}(w)\|\leq\mathtt{e}$. Within our final numerical implementation in Section 5.3, we allow for nonexpansive operators, corresponding to $q=1$, heavily noting that the theoretical guarantees are weakened in this setting.
Remark 3.3.
Rather than assuming exact access to the operators $\mathsf{T}_{k}$, we could assume access to $\delta_{k}$-perturbations $\tilde{\mathsf{T}}_{k}$, with bounded inexactness model $0\leq\delta_{k}\leq\delta_{\max}$. Such an error structure is motivated by the fact that the exact map might in general not be available, or its computation may be very demanding. This occurs for instance when applying proximal methods to image deblurring with total variation [Chambolle:2004aa], or to structured sparsity regularization problems in machine learning and inverse problems [Zhao:2009aa]. In those cases, the proximity operator is usually computed using ad hoc algorithms, and therefore inexactly. The implications of inexact implementations are the following: Lemma 2.1 would still hold, by replacing $\varepsilon$ by $\max(\varepsilon,\delta_{\max})$, provided $\delta_{\max}$ is finite. Additionally, Theorem 3.1 would still hold [Cortild:2024aa], provided $(\theta_{k}\delta_{k})\in\ell^{2}$, but without an explicit rate. As such, Corollary 3.2 still holds, but again without rate guarantees. Specifically, we have no apriori upper bound on the required number of iterations required for Algorithm 1 to terminate, but we know it will terminate.
3.1.1 Parameters Analysis
We want to take a closer look at the numerical parameters that can be chosen in the inner loop. We restrict ourselves to the case of constant parameters, namely $\tau_{k}\equiv\tau\in(0,1)$, $\theta_{k}\equiv\theta\in(0,1)$, and $q_{k}\equiv q\in(0,1)$, implied by $\mathsf{T}_{k}\equiv\mathsf{T}$. In this case, Inequality (3.1) in Assumption 3 becomes
$\tau^{2}\cdot(1-2\theta+\theta^{2}(1-q^{2}))+\tau\cdot(2-(2-q^{2})\theta)-(1-(1-q^{2})\theta)(1-\theta)<0.$
As such, for a fixed $(\theta,q)$, the admissible values of $\tau$ are $\tau\in[0,\overline{\tau})$, where $\overline{\tau}$ is the largest root of the given quadratic. In Figure 1, we plot $\overline{\tau}$ for different values of $q$.
Figure 1: Largest acceleration parameter $\overline{\tau}$ as a function of the relaxation and the contraction parameters $\theta$ and $q$.
3.2 Convergence Analysis of Restarting Procedure
In the following, we assume that the operator $\mathsf{G}$ is $\mu$-strongly monotone with $\mu\geq 0$. The case in which $\mathsf{G}$ is merely monotone is recovered by setting $\mu=0$ throughout the subsequent analysis.
We introduce two gap functions, one associated with optimality and one with feasibility.
The optimality gap is defined as
$\operatorname{Gap}_{\operatorname{opt}}(u)\triangleq\sup_{v\in\mathcal{S}_{0}}\langle\mathsf{G}v,u-v\rangle.$ ($\operatorname{Gap}_{\operatorname{opt}}$)
We observe that if $u$ is feasible, namely if $u\in\mathcal{S}_{0}=\operatorname{Zer}(\mathsf{M})$, then the condition $\operatorname{Gap}_{\operatorname{opt}}(u)\leq 0$ is equivalent to $u\in\operatorname{Zer}(\mathsf{G}+\operatorname{\mathsf{NC}}_{\mathcal{S}_{0}})$, such that $u$ solves Problem (P). It is important to note, however, that $\operatorname{Gap}_{\operatorname{opt}}(u)\leq 0$ provides limited information when $u\notin\operatorname{Zer}(\mathsf{M})$. It is thus important to obtain a lower bound on the optimality gap.
The feasibility gap is motivated by the observation that, by maximal monotonicity of $\mathsf{M}$, for any $x\in\mathcal{H}$,
$x\in\mathcal{S}_{0}=\operatorname{Zer}(\mathsf{M})\Leftrightarrow\inf_{y\in\operatorname{dom}(\mathsf{M})}\inf_{v\in\mathsf{M}y}\langle 0-v,x-y\rangle\geq 0.$
As such, we define the feasibility gap $\operatorname{Gap}_{\operatorname{feas}}:\mathcal{H}\to\mathbb{R}$ as
$\operatorname{Gap}_{\operatorname{feas}}(x)=\sup_{y\in\operatorname{dom}(\mathsf{M}),v\in\mathsf{M}y}\langle v,x-y\rangle.$ ($\operatorname{Gap}_{\operatorname{feas}}$)
We note that $\operatorname{Gap}_{\operatorname{feas}}(x)\leq 0$ implies $x\in\operatorname{Zer}(\mathsf{M})$, and that $\operatorname{Gap}_{\operatorname{feas}}(x)\geq 0$ holds for all $x\in\mathcal{H}$ since, by assumption, the problem admits at least one solution; in particular, $\operatorname{Zer}(\mathsf{M})\neq\varnothing$. Since $\operatorname{Gap}_{\operatorname{opt}}$ and $\operatorname{Gap}_{\operatorname{feas}}$ are pointwise suprema of affine functions, both are weakly lower semicontinuous.
Our convergence rates in terms of the optimality gap require a lower bound on the a priori signless optimality gap function $\operatorname{Gap}_{\operatorname{opt}}$. We derive such a lower bound within a specific geometric setting, imposing a Hölderian error bound on the feasibility gap.
Definition 3.4.
We say that the zero set $\mathcal{S}_{0}=\operatorname{Zer}(\mathsf{M})$ is weakly sharp with constant $\kappa>0$ and order $\rho\geq 1$ if, for all $u\in\operatorname{dom}(\mathsf{M})$,
$\kappa\cdot\operatorname{dist}(u,\operatorname{Zer}(\mathsf{M}))^{\rho}\leq\operatorname{Gap}_{\operatorname{feas}}(u).$ (3.2)
The notion of weak sharpness of zero sets is closely related to the classical concept of weak sharpness for variational inequalities [marcotte_weak_1998]. Geometric conditions ensuring the error-bound characterization (3.2) can be found in [Al-Homidan:2017aa]. The recent papers [samadi_improved_2025] and [Alvarez:2001aa] rely also heavily on this assumption. An important class of examples arises when $\mathsf{M}$ is single-valued and $\rho=1$, notably in monotone linear complementarity problems in finite dimensions under nondegeneracy conditions [burke_weak_1993, pang1997error]. In the case where the lower-level problem is determined by a convex optimization problem, weak-sharpness implies an Hölderian errror bound, an assumption already imposed by [Cabot2005] in the context of hierarchical minimization. Specifically, let us assume that $\mathsf{M}=\partial f$ for a proper convex and lower semi-continuous function $f:\mathcal{H}\to\mathbb{R}\cup\{+\infty\}$. Then,
$\sup_{v\in\partial f(y)}\langle v,u-y\rangle=f^{\prime}(y;u-y)\leq f(u)-f(y),$
so that $\operatorname{Gap}_{\operatorname{feas}}(u)\leq f(u)-\min f$. It follows that Inequality (3.2) implies
$\kappa\operatorname{dist}(u,\operatorname*{argmin}f)^{\rho}\leq f(u)-\min f\hskip 17.00024pt\forall u\in\mathcal{H}.$
If $\rho=1$, this is the weak-sharpness condition of [burke_weak_1993]. The case $\rho=2$ corresponds to the "quadratic growth" condition of [drusvyatskiy2018error]. Rates of convergence under this Hölderian error bound assumption have recently been derived in [boct2025accelerating].
We next define the following quantities, which are finite because $\operatorname{dom}(\mathsf{M})$ is assumed to be bounded:
$D_{\mathsf{M}}=\operatorname{diam}(\operatorname{dom}(\mathsf{M})),\hskip 17.00024pt\hskip 17.00024pt$ $C_{\mathsf{M}}=\sup_{x\in\operatorname{dom}(\mathsf{M})}\sup_{v\in\mathsf{M}x}\|v\|,\qquad$ $C_{\mathsf{G}}=\sup_{x\in\operatorname{dom}(\mathsf{M})}\mathsf{G}x.$
We also give more explicit expressions for the averaging coefficients $(\lambda_{n})_{n}$ and the corresponding averaged iterates which are produced by DANTE. They read as
$\lambda_{0}=1,\text{ and for all }n\geq 1,\;\lambda_{n}=\prod_{i=0}^{n-1}\left(1+\frac{2\mu\beta_{i}}{\alpha}\right),\;\overline{w}_{N}=\frac{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}w_{n+1}}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}.$
In order to guarantee our convergence results, we must make some weak assumptions on the sequence of parameters.
Assumption 5.
We assume the following on the parameter sequences:
  1. 1.
    $\beta_{n}\in[0,\beta_{\max}]$ is non-increasing and $\lim_{n\to\infty}\beta_{n}=0$.
  2. 2.
    The sequence $(\lambda_{n}\beta_{n})_{n\geq 0}$ satisfies $\sum_{n\geq 0}\lambda_{n}\beta_{n}=\infty$, and
    $\frac{\sum_{n=0}^{N-1}\lambda_{n}\beta^{2}_{n}}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}\to 0\text{ as }N\to\infty.$ (3.3)
  3. 3.
    The sequence $(\mathtt{e}_{n})_{n\geq 0}$ satisfies $\mathtt{e}_{n}=o(\beta_{n})$. We define $\mathtt{e}_{\max}=\max_{n\geq 1}\mathtt{e}_{n}$.
Note that since $\mathtt{e}_{n}=O(\varepsilon_{n})$, Assumption 5 implies $\varepsilon_{n}/\beta_{n}\to 0$ as $n\to\infty.$ This means that when designing concrete parameters for implementing our algorithm, we can tailor the inner-loop tolerance $\varepsilon_{n}$ to the evolution of the Tikhonov parameter $\beta_{n}\to 0$.
We are now ready to state the proposition that is central to our analysis.
Proposition 3.5.
For any $x\in\operatorname{dom}(\mathsf{M})$ and $v\in\mathsf{M}x$, it holds that
$\lambda_{n}\langle v,w_{n+1}-x\rangle+\lambda_{n}\beta_{n}\langle\mathsf{G}x,w_{n+1}-x\rangle\leq-\frac{\alpha\lambda_{n+1}}{2}\|w_{n+1}-x\|^{2}+\frac{\alpha\lambda_{n}}{2}\|w_{n}-x\|^{2}+C_{1}\lambda_{n}\mathtt{e}_{n},$ (3.4)
where
$C_{1}\triangleq 2\alpha\left(1+\frac{\mu\beta_{\max}}{\alpha}\right)(\mathtt{e}_{\max}+D_{\mathsf{M}})+\frac{C_{\mathsf{M}}}{\alpha}+\frac{\beta_{\max}C_{\mathsf{G}}}{\alpha}.$ (3.5)
Proof.
First, expand
$\frac{1}{2}\|w_{n+1}-x\|^{2}-\frac{1}{2}\|w_{n}-x\|^{2}$ $=-\frac{1}{2}\|w_{n+1}-w_{n}\|^{2}+\langle w_{n+1}-w_{n},w_{n+1}-x\rangle$
$\leq\langle w_{n}-\overline{u}_{n+1},x-\overline{u}_{n+1}\rangle+\langle w_{n+1}-w_{n},w_{n+1}-\overline{u}_{n+1}\rangle$
$\qquad+\langle\overline{u}_{n+1}-w_{n+1},x-\overline{u}_{n+1}\rangle.$
For the last two terms, we obtain the bounds
$\langle w_{n+1}-w_{n},w_{n+1}-\overline{u}_{n+1}\rangle$ $\leq\left(\|{w_{n+1}-\overline{u}_{n+1}}\|+\|{\overline{u}_{n+1}-\overline{u}_{n}}\|+\|{\overline{u}_{n}-w_{n}}\|\right)\cdot\|{w_{n+1}-\overline{u}_{n+1}}\|$
$\leq(\mathtt{e}_{n}+D_{\mathsf{M}}+\mathtt{e}_{n-1})\mathtt{e}_{n}$
$\leq(D_{\mathsf{M}}+2\mathtt{e}_{max})\mathtt{e}_{n},$
and
$\langle\overline{u}_{n+1}-w_{n+1},\overline{u}_{n+1}-x\rangle\leq\|{\overline{u}_{n+1}-w_{n+1}}\|\cdot\|{\overline{u}_{n+1}-x}\|\leq D_{\mathsf{M}}\mathtt{e}_{n}.$
Combining these inequalities yields
$\frac{1}{2}\|w_{n+1}-x\|^{2}-\frac{1}{2}\|w_{n}-x\|^{2}\leq\langle w_{n}-\overline{u}_{n+1},x-\overline{u}_{n+1}\rangle+2(D_{\mathsf{M}}+\mathtt{e}_{\max})\mathtt{e}_{n}.$
As $\beta_{n}\mathsf{G}\overline{u}_{n+1}-\alpha(\overline{u}_{n+1}-w_{n})\in\mathsf{M}\overline{u}_{n+1}$ and $v\in\mathsf{M}x$, it follows by monotonicity of $\mathsf{M}$ and by $\mu$-monotonicity of $\mathsf{G}$ that
$\langle w_{n}-\overline{u}_{n+1},x-\overline{u}_{n+1}\rangle$ $\leq\frac{1}{\alpha}\langle v,x-\overline{u}_{n+1}\rangle+\frac{\beta_{n}}{\alpha}\langle\mathsf{G}\overline{u}_{n+1},x-\overline{u}_{n+1}\rangle$
$\leq\frac{1}{\alpha}\langle v,x-\overline{u}_{n+1}\rangle+\frac{\beta_{n}}{\alpha}\langle\mathsf{G}x,x-\overline{u}_{n+1}\rangle-\frac{\mu\beta_{n}}{\alpha}\|x-\overline{u}_{n+1}\|^{2}.$
Now we note that
$\|x-\overline{u}_{n+1}\|^{2}$ $=\|x-w_{n+1}\|^{2}+\|w_{n+1}-\overline{u}_{n+1}\|^{2}+2\langle x-w_{n+1},w_{n+1}-\overline{u}_{n+1}\rangle$
$\geq\|x-w_{n+1}\|^{2}-2(\|x-\overline{u}_{n+1}\|+\|\overline{u}_{n+1}-w_{n+1}\|)\cdot\|w_{n+1}-\overline{u}_{n+1}\|$
$\geq\|x-w_{n+1}\|^{2}-2(D_{\mathsf{M}}+\mathtt{e}_{\max})\mathtt{e}_{n}.$
Combining the previous three relationships yields
$\frac{1}{2}\|w_{n+1}-x\|^{2}-\frac{1}{2}\|w_{n}-x\|^{2}$ $\leq\frac{1}{\alpha}\langle v,x-\overline{u}_{n+1}\rangle+\frac{\beta_{n}}{\alpha}\langle\mathsf{G}x,x-\overline{u}_{n+1}\rangle-\frac{\mu\beta_{n}}{\alpha}\|{x-w_{n+1}}\|^{2}$
$\qquad+2\frac{\mu\beta_{n}}{\alpha}(D_{\mathsf{M}}+\mathtt{e}_{max})\mathtt{e}_{n}+2(D_{\mathsf{M}}+\mathtt{e}_{\max})\mathtt{e}_{n}.$
Rearranging the terms and bounding $\beta_{n}\leq\beta_{\max}$ gives
$\left(\frac{1}{2}+\frac{\mu\beta_{n}}{\alpha}\right)\|w_{n+1}-x\|^{2}-\frac{1}{2}\|w_{n}-x\|^{2}\leq\frac{1}{\alpha}\langle v,x-\overline{u}_{n+1}\rangle+\frac{\beta_{n}}{\alpha}\langle\mathsf{G}x,x-\overline{u}_{n+1}\rangle+2\left(1+\frac{\mu\beta_{\max}}{\alpha}\right)(D_{\mathsf{M}}+\mathtt{e}_{\max})\mathtt{e}_{n}.$
After noting that
$\langle v,x-\overline{u}_{n+1}\rangle=\langle v,x-w_{n+1}\rangle+\langle v,w_{n+1}-\overline{u}_{n+1}\rangle\leq\langle v,x-w_{n+1}\rangle+C_{\mathsf{M}}\mathtt{e}_{n},$
and
$\langle\mathsf{G}x,x-\overline{u}_{n+1}\rangle=\langle\mathsf{G}x,x-w_{n+1}\rangle+\langle\mathsf{G}x,w_{n+1}-\overline{u}_{n+1}\rangle\leq\langle\mathsf{G}x,x-w_{n+1}\rangle+C_{\mathsf{G}}\mathtt{e}_{n},$
and multiplying both sides by $\lambda_{n}\alpha$, we obtain
$\lambda_{n}\langle v,w_{n+1}-x\rangle$ $+\lambda_{n}\beta_{n}\langle\mathsf{G}x,w_{n+1}-x\rangle$
$\leq-\left(\frac{1}{2}+\frac{\mu\beta_{n}}{\alpha}\right)\lambda_{n}\alpha\|w_{n+1}-x\|^{2}+\frac{\lambda_{n}\alpha}{2}\|{w_{n}-x}\|^{2}$
$\qquad+\left(C_{\mathsf{M}}+\beta_{max}C_{\mathsf{G}}+2\left(1+\frac{\mu\beta_{\max}}{\alpha}\right)(\mathtt{e}_{\max}+D_{\mathsf{M}})\alpha\right)\lambda_{n}\mathtt{e}_{n}.$
The results follow by recalling the definition of $\lambda_{n}$. $\blacksquare$
The lower bound in the preceding proposition enables us to deduce bounds on both the optimality and feasibility gaps, provided the remaining term can be controlled. The next two lemmas accomplish precisely this.
Lemma 3.6 (Optimality).
For all $N\geq 1$, it holds that
$\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{N})\leq C_{2}\cdot\frac{1}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}+C_{1}\cdot\frac{\sum_{n=0}^{N-1}\lambda_{n}\mathtt{e}_{n}}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}},$
where $C_{1}$ is defined in eq. (3.5) and
$C_{2}\triangleq\alpha(\mathtt{e}_{0}^{2}+D_{\mathsf{M}}^{2}).$ (3.6)
In particular, if Assumption 5 holds, all weak accumulation points $\tilde{w}$ of $(\overline{w}_{n})$ satisfy $\operatorname{Gap}_{\operatorname{opt}}(\tilde{w})\leq 0$.
Proof.
Select $x\in\operatorname{Zer}(\mathsf{M})$ and $v=0$ in (3.4), and sum the result for $n=0,\ldots,N-1$ to obtain that
$\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}\langle\mathsf{G}x,w_{n+1}-x\rangle$ $\leq\frac{-\alpha\lambda_{N}}{2}\|w_{N}-x\|^{2}+\frac{\alpha\lambda_{0}}{2}\|w_{0}-x\|^{2}+C_{1}\cdot\sum_{n=0}^{N-1}\lambda_{n}\mathtt{e}_{n}$
$\leq{\alpha\lambda_{0}}(\|w_{0}-\overline{u}_{0}\|^{2}+\|\overline{u}_{0}-x\|^{2})+C_{1}\cdot\sum_{n=0}^{N-1}\lambda_{n}\mathtt{e}_{n}$
$\leq{\alpha}(\mathtt{e}_{0}^{2}+D_{\mathsf{M}}^{2})+C_{1}\cdot\sum_{n=0}^{N-1}\lambda_{n}\mathtt{e}_{n}.$
After normalization, this yields
$\langle\mathsf{G}x,\overline{w}_{N}-x\rangle\leq C_{2}\cdot\frac{1}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}+C_{1}\cdot\frac{\sum_{n=0}^{N-1}\lambda_{n}\mathtt{e}_{n}}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}.$
The inequality follows by taking the supremum over $x\in\operatorname{Zer}(\mathsf{M})$. The convergence conclusion follows by weak sequential lower-semicontinuity of $\operatorname{Gap}_{\operatorname{opt}}$. $\blacksquare$
Lemma 3.7 (Feasibility).
Assume $(\beta_{n})$ non-increasing. For all $N\geq 1$, it holds that
$0\leq\operatorname{Gap}_{\operatorname{feas}}(\overline{w}_{N})\leq C_{3}\cdot\frac{1}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}+C_{4}\cdot\frac{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}^{2}}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}+C_{1}\cdot\frac{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}\mathtt{e}_{n}}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}},$
where
$C_{3}=\beta_{0}C_{2},\qquad C_{4}\triangleq C_{\mathsf{G}}(\mathtt{e}_{\max}+D_{\mathsf{M}})$ (3.7)
where $C_{1}$ and $C_{2}$ are defined in Equations (3.5) and (3.6). In particular, if Assumption 5 holds, all weak cluster point $\tilde{w}$ of $(\overline{w}_{n})$ satisfy $\operatorname{Gap}_{\operatorname{feas}}(\tilde{w})=0$, such that $\tilde{w}\in\operatorname{Zer}(\mathsf{M})$.
Proof.
Multiply both sides of (3.4) by $\beta_{n}$, and recall that $\beta_{n+1}\leq\beta_{n}$. Moreover, notice that
$\langle\mathsf{G}x,w_{n+1}-x\rangle\geq-\|{\mathsf{G}x}\|\cdot\|{w_{n+1}-x}\|\geq-\|\mathsf{G}x\|\cdot(\|w_{n+1}-\overline{u}_{n+1}\|+\|\overline{u}_{n+1}-x\|)\geq-C_{\mathsf{G}}(\mathtt{e}_{n}+D_{\mathsf{M}}).$
It then follows $\langle\mathsf{G}x,w_{n+1}-x\rangle\leq C_{\mathsf{G}}(\mathtt{e}_{\max}+D_{\mathsf{M}})$. Summing for $n=0,\ldots,N-1$ then yields
$\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}\langle v,w_{n+1}-x\rangle$ $\leq-\frac{\alpha\lambda_{N}\beta_{N}}{2}\|w_{N}-x\|^{2}+\frac{\alpha\lambda_{0}\beta_{0}}{2}\|w_{0}-x\|^{2}+C_{1}\cdot\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}\mathtt{e}_{n}$
$\qquad+C_{\mathsf{G}}(\mathtt{e}_{\max}+D_{\mathsf{M}})\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}^{2}$
$\leq C_{3}+C_{1}\cdot\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}\mathtt{e}_{n}+C_{4}\cdot\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}^{2}.$
By dividing by $\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}$ and taking the supremum over points $x\in\mathcal{S}_{0}$ on both sides, we obtain the claimed inequality. The convergence conclusion again follows by weak sequential lower semicontinuity of $\operatorname{Gap}_{\operatorname{feas}}$. $\blacksquare$
We may now combine the two previous lemmas to state our main theorem.
Theorem 3.8 (Main Theorem).
Let $(w_{n})_{n=0}^{N-1}$ be generated through Algorithm 2. Assume $(\beta_{n})_{n\geq 0}$ is non-increasing. Then the following statements hold true:
  • (a)
    For all $N\geq 1$, the optimality gap is bounded as
    $C_{\mathsf{G}}\cdot\operatorname{dist}(\overline{w}_{N},\mathcal{S}_{0})\leq\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{N})\leq C_{2}\cdot\frac{1}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}+C_{1}\cdot\frac{\sum_{n=0}^{N-1}\lambda_{n}\mathtt{e}_{n}}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}},$ (3.8)
    where $C_{1}$ and $C_{2}$ are problem-specific constants defined in (3.5) and (3.6).
  • (b)
    For all $N\geq 1$, the feasibility gap is bounded as
    $0\leq\operatorname{Gap}_{\operatorname{feas}}(\overline{w}_{N})\leq C_{3}\cdot\frac{1}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}+C_{4}\cdot\frac{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}^{2}}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}+C_{1}\cdot\frac{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}\mathtt{e}_{n}}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}},$ (3.9)
    where $C_{3}$ and $C_{4}$ are constants defined in (3.7).
  • (c)
    With Assumption 5 in place, all weak accumulation points of $(\overline{w}_{n})_{n\geq 1}$ are solutions to (P).
  • (d)
    If, moreover, $\mathcal{S}_{0}$ is weakly sharp with constant $\kappa$ or order $\rho\geq 1$, then it also holds that
    $\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{N})\geq-C_{\mathsf{G}}\kappa^{-1/\rho}\cdot\left(C_{3}\cdot\frac{1}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}+C_{4}\cdot\frac{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}^{2}}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}+C_{1}\cdot\frac{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}\mathtt{e}_{n}}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}\right)^{1/\rho}.$
Proof.
  • (a)
    The upper bound is just Lemma 3.6. For the lower bound, we invoke the Cauchy-Schwarz inequality $\langle\mathsf{G}x,\overline{w}_{N}-x\rangle\geq\|{\mathsf{G}x}\|\cdot\|{\overline{w}_{N}-x}\|$, implying
    $\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{N})\geq-C_{\mathsf{G}}\operatorname{dist}(\overline{w}_{N},\mathcal{S}_{0}).$
    This proves the assertion.
  • (b)
    This is just Lemma 3.7.
  • (c)
    Combine Lemma 3.6 with Lemma 3.7.
  • (d)
    If $\mathcal{S}_{0}$ is weakly sharp with constant $\kappa$ of order $\rho\geq 1$, then
    $\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{N})$ $\geq-C_{\mathsf{G}}\operatorname{dist}(\overline{w}_{N},\mathcal{S}_{0})$
    $\geq-C_{\mathsf{G}}\left[\frac{1}{\kappa}\operatorname{Gap}_{\operatorname{feas}}(\overline{w}_{n})\right]^{1/\rho}$
    $\geq-\frac{C_{\mathsf{G}}}{\kappa^{1/\rho}}\left(C_{3}\cdot\frac{1}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}+C_{4}\cdot\frac{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}^{2}}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}+C_{1}\cdot\frac{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}\mathtt{e}_{n}}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}\right)^{1/\rho}.$
$\blacksquare$
By selecting specific regularization parameters, we may obtain explicit rates on the convergent quantities, allowing direct comparisons to prior works. The choice of parameters differs between the solely monotone setting, detailed below, and the strongly monotone setting, detailed in Corollary 3.10.
Corollary 3.9.
Let $(w_{n})$ be generated through Algorithm 2 when $\mathsf{G}$ is merely monotone (i.e. $\mu=0$). Let $\beta_{n}=(n+1)^{-b}$ for $b\in(0,1)$ and choose $\varepsilon_{n}$ such that $\mathtt{e}_{n}=o(\beta_{n})$. Then, for all $N\geq 2^{1/(1-b)}$, it holds that
$\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{N})\leq\mathcal{O}\left(\frac{1}{N^{1-b}}\right),$
as well
$0\leq\operatorname{Gap}_{\operatorname{feas}}(\overline{w}_{N})\leq\begin{cases}\mathcal{O}\left(\frac{1}{N^{b}}\right)&\text{if $b\in(0,0.5)$}\\ \mathcal{O}\left(\frac{\ln(N)}{N^{1/2}}\right)&\text{if $b=0.5$}\\ \mathcal{O}\left(\frac{1}{N^{1-b}}\right)&\text{if $b\in(0.5,1)$}.\end{cases}$
If, moreover, $\mathcal{S}_{0}$ is weakly sharp with constant $\kappa>0$ of order $\rho\geq 1$, then it also holds that
$\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{N})\geq\mathcal{O}(\operatorname{Gap}_{\operatorname{feas}}(\overline{w}_{N})^{1/\rho}).$
Proof.
By standard integral bounds, one can show that for $c\in(0,1)$ and $N\geq 2^{1/(1-b)}$,
$\frac{N^{1-c}}{2(1-c)}\leq\sum_{n=0}^{N-1}(n+1)^{-c}\leq\frac{N^{1-c}}{1-c}.$
Specifically, $\sum_{n=0}^{N-1}\beta_{n}=\Theta(N^{1-b})$, and
$\frac{\sum_{n=0}^{N-1}\beta_{n}^{2}}{\sum_{n=0}^{N-1}\beta_{n}}=\begin{cases}\Theta(N^{-b})&\text{if $b\in(0,0.5)$}\\ \Theta(\ln(N)N^{-1/2})&\text{if $b=0.5$}\\ \Theta(N^{b-1})&\text{if $b\in(0.5,1)$}\end{cases}.$
As $\mathtt{e}_{n}=o(\beta_{n})$, the conclusions follow by Theorem 3.8. $\blacksquare$
Corollary 3.10.
Let $(w_{n})_{n=0}^{N}$ be generated through Algorithm 2 when $\mathsf{G}$ is strongly monotone (i.e. $\mu>0$). Let $\beta_{n}=\frac{\alpha}{2\mu n+\xi}$ for $\xi>0$ and let $\mathtt{e}_{n}=o(\beta_{n})$. For all $N$ sufficiently large, it holds that
$\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{N})\leq\mathcal{O}\left(\frac{1}{N}\right),$
and that
$0\leq\operatorname{Gap}_{\operatorname{feas}}(\overline{w}_{N})\leq\mathcal{O}\left(\frac{\log(N)}{N}\right).$
If, moreover, $\operatorname{Zer}(\mathsf{M})$ is weakly sharp with constant $\kappa$ of order $\rho\geq 1$, then it also holds that $\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{N})\geq\mathcal{O}(\operatorname{Gap}_{\operatorname{feas}}(\overline{w}_{N})^{1/\rho}).$
Proof.
Under the stated hypothesis, one easily computes that $\lambda_{n}=\frac{2\mu n+\xi}{\xi}$. As such,
$\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}=\frac{N\alpha}{\xi},$
and
$\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}^{2}=\frac{\alpha^{2}}{\xi}\cdot\sum_{n=0}^{N-1}\frac{1}{2\mu n+\xi}\leq\frac{\alpha^{2}}{\xi}\cdot\left(\frac{1}{\xi}+\frac{1}{2\mu}\log\frac{2\mu N+\xi}{\xi}\right)=\mathcal{O}(\log(N)).$
The conclusion follows by applying these estimates to Theorem 3.8. $\blacksquare$
3.3 Combined Complexity Analysis
We now finally perform the complete complexity analysis of our scheme DANTE by combining our bounds derived for Algorithms 1 and 2.
Before stating the final convergence result, we make one final assumption.
Assumption 6.
We assume $\mathsf{Z}_{n}$ is domain-forward on $\operatorname{dom}(\mathsf{M})$, namely that $\mathsf{Z}_{n}(\operatorname{dom}(\mathsf{M}))\subset\operatorname{dom}(\mathsf{M})$. Finally, we assume that $\theta_{n,k}$ is uniformly lower bounded by a positive quantity, such that $\overline{Q}_{n}\leq\overline{Q}<1$.
Remark 3.4.
The domain-forwardness assumption may appear restrictive on the first glance. However, for all relevant fixed-point mappings from Section 4, the fixed-point transportation map will take the form of a resolvent operator for which the requirement is naturally fulfilled due to maximal monotonicity of the operator.
Theorem 3.11.
Assume the setting of Corollary 3.3 and that Assumptions 2, 3, 4 and 6 hold. Algorithm 2 (DANTE) reaches a point $\overline{w}_{N}$ satisfying the inequalities in Theorem 3.8 after at most
$K=\frac{N(2\log(C)+\log(\overline{Q}))+2\sum_{n=1}^{N}\log(1/\varepsilon_{n})}{\log(1/\overline{Q})}$
inner iterations. If $\varepsilon_{n}=(n+1)^{-\eta}$ where $\eta\geq b$, we reach the desired point $\overline{w}_{N}$ after at most ${\mathcal{O}}(N\log N)$ inner iterations.
Proof.
By Corollary 3.3, each outer loop $n\in\{0,\ldots,N-1\}$ runs at most
$\overline{K}_{n}=\left\lceil\frac{2\log(C)+2\log(1/\varepsilon_{n})}{\log(1/\overline{Q})}\right\rceil\leq\frac{2\log(C)+2\log(1/\varepsilon_{n})}{\log(1/\overline{Q})}+1$
inner iterations. Specifically, the inequalities of Theorem 3.8 are satisfied after at most
$K=\sum_{n=0}^{N-1}\overline{K}_{n}\leq\frac{2N\log(C)+2\sum_{n=1}^{N}\log(1/\varepsilon_{n})}{\log(1/\overline{Q})}+N=\mathcal{O}\left(N+\sum_{n=0}^{N-1}\log(1/\varepsilon_{n})\right)$
By setting $\varepsilon_{n}=(n+1)^{-\eta}$, we obtain that $\sum_{n=0}^{N-1}\log(1/\varepsilon_{n})\leq\eta N\log(N)$, as wanted. $\blacksquare$
Remark 3.5.
Theorem 3.11 shows that accounting for the total number of inner iterations, rather than only the number of outer iterations, introduces merely a logarithmic factor. This is expected, as double-loop schemes inherently incur such a term. Consequently, the complexities stated in Corollaries 3.9 and 3.10 are preserved, up to logarithmic factors, when measured in terms of total inner iterations.
4 Relevant Splitting Methods
Depending on the structure of $\mathsf{M}$ in Problem (P), the auxiliary problem (Aux) inherits different structural assumptions. The fixed-point encodings $\mathsf{T}^{(w,\beta)}_{k}$ must be readily available for the numerical computations, and hence should exploit known problem structure. In Section 4.1, we analyze the case where $\mathsf{M}$ may be written as the sum of two operators, and in Section 4.2, we consider a three-operator splitting scheme.
Remark 4.1.
With the following examples it will become clear that the fixed-point encoding $\mathsf{T}^{(w,\beta)}_{k}$ not only depends on the problem parameters $\alpha$ and $\beta$ but also on the step size $\gamma_{k}$. Therefore, we treat the step size policy is treated in our approach as an ingredient of the fixed-point encoding strategy and consequently incorporated in the map $\mathsf{T}^{(w,\beta)}_{k}$, rather than being a parameter of the inner loop presented in Algorithm 1.
4.1 Two-Operator Splitting Schemes
In this section, we focus on problems of the form (P) with $\mathsf{M}=\mathsf{A}+\mathsf{F}$, namely
$\operatorname{VI}(\mathsf{G},\mathcal{S}_{0}),\qquad\text{where }\mathcal{S}_{0}\triangleq\operatorname{Zer}(\mathsf{A}+\mathsf{F}).$ (P-2-Split)
We refine the problem assumptions through the following assumption:
Assumption 7.
We assume the following on Problem (P-2-Split):
  1. 1.
    The lower level solution set $\mathcal{S}_{0}$ is nonempty;
  2. 2.
    $\mathsf{G}:\mathcal{H}\to\mathcal{H}$ is monotone and $L_{\mathsf{G}}$-Lipschitz continuous;
  3. 3.
    $\mathsf{F}\colon\mathcal{H}\to\mathcal{H}$ is monotone and $L_{\mathsf{F}}$-Lipschitz continuous;
  4. 4.
    $\mathsf{A}\colon\mathcal{H}\to 2^{\mathcal{H}}$ is maximally monotone with bounded domain.
These assumptions imply that $\mathsf{M}=\mathsf{A}+\mathsf{F}$ is maximally monotone with $\operatorname{dom}(\mathsf{M})=\operatorname{dom}(\mathsf{A})$.
Given the parameters $\alpha$ and $\beta$ we define the function $\Phi^{w}_{\alpha,\beta}:\mathcal{H}\to\mathcal{H}$ by
$\Phi^{w}_{\alpha,\beta}(v)\triangleq\mathsf{F}(v)+\beta\mathsf{G}(v)+\alpha(v-w),$
which is Lipschitz continuous with constant $L_{\alpha,\beta}\triangleq L_{\mathsf{F}}+\beta L_{\mathsf{G}}+\alpha$ and $\alpha$-strongly monotone. Adopting this notation, we can rewrite the auxiliary problem (Aux) as determining the unique element of $\operatorname{Zer}(\mathsf{A}+\Phi_{\alpha,\beta}^{w})$.
Forward-backward splitting.
The Forward-Backward splitting [lions_splitting_1979, passty_ergodic_1979] is a very popular numerical scheme, defined in terms of the fixed-point encoding map $\mathsf{T}^{(w,\beta)}_{\gamma}\colon\mathcal{H}\to\mathcal{H}$ given by
$\mathsf{T}^{(w,\beta)}_{\gamma}=\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}\circ(\operatorname{Id}-\gamma\Phi_{\alpha,\beta}^{w}).$ (FB)
We note that $\operatorname{Fix}(\mathsf{T}^{(w,\beta)}_{\gamma})=\operatorname{Zer}(A+\Phi_{\alpha,\beta}^{w})$, so $\mathsf{Z}=\operatorname{Id}$ is the fixed-point transportation map. This is the setting studied in [marschner_tikhonov_2025].
Lemma 4.1.
Assume Assumption 7 holds and let $\mathsf{T}^{(w,\beta)}_{\gamma}$ be defined as (FB). $\mathsf{T}^{(w,\beta)}_{\gamma}$ is $q$-Lipschitz where
$q\triangleq\sqrt{1-\gamma(2\alpha-\gamma L^{2}_{\alpha,\beta})},$
uniformly over $w\in\mathcal{H}$. Hence, $\mathsf{T}^{(w,\beta)}_{\gamma}$ is a contraction when $\gamma\in(0,2\alpha/L^{2}_{\alpha,\beta})$.
Proof.
Pick $z_{1},z_{2}\in\mathcal{H}$. By nonexpansiveness of $\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}$, and $L_{\alpha,\beta}$-Lipschitz continuity and $\alpha$-strong monotonicity of $\Phi_{\alpha,\beta}$,
$\|{T(z_{1})-T(z_{2})}\|^{2}$ $=\|{\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}(z_{1}-\gamma\Phi_{\alpha,\beta}(z_{1},w))-\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}(z_{2}-\gamma\Phi_{\alpha,\beta}(z_{2},w))}\|^{2}$
$\leq\|{z_{1}-z_{2}+\gamma(\Phi_{\alpha,\beta}(z_{2},w)-\Phi_{\alpha,\beta}(z_{2},w))}\|^{2}$
$=\|{z_{1}-z_{2}}\|^{2}+2\langle z_{1}-z_{2},\Phi_{\alpha,\beta}(z_{2},w)-\Phi_{\alpha,\beta}(z_{1},w)\rangle$
$\hskip 17.00024pt+\gamma^{2}\|{\Phi_{\alpha,\beta}(z_{1},w)-\Phi_{\alpha,\beta}(z_{2},w)}\|^{2}$
$\leq\|{z_{1}-z_{2}}\|^{2}-2\gamma\alpha\|{z_{1}-z_{2}}\|^{2}+L^{2}_{\alpha,\beta}\gamma^{2}\|{z_{1}-z_{2}}\|^{2}$
$=(1-2\gamma\alpha+L^{2}_{\alpha,\beta}\gamma^{2})\cdot\|{z_{1}-z_{2}}\|^{2}.$
The contraction property follows by choosing $\gamma$ so that $\gamma(2\alpha-\gamma L^{2}_{\alpha,\beta})\in(0,1)$. $\blacksquare$
Backward-forward splitting.
In a similar way, we can employ the Backward-Forward splitting [attouch_backward_2018] for the auxiliary problem. Consider the operator
$\mathsf{T}^{(w,\beta)}_{\gamma}=(\operatorname{Id}-\gamma\Phi_{\alpha,\beta}^{w})\circ\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}.$ (BF)
Note that $\operatorname{Zer}(\mathsf{A}+\Phi_{\alpha,\beta}^{w})=\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}(\text{Fix}(\mathsf{T}^{(w,\beta)}_{\gamma}))$, which means that the last iterate of the inner loop scheme needs one more evaluation of the resolvent of the operator $\mathsf{A}$ in order to get close to the unique zero of $\mathsf{A}+\Phi_{\alpha,\beta}^{w}$. Specifically, the fixed-point transportation map is $\mathsf{Z}=\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}$.
Lemma 4.2.
Assume Assumption 7 holds and let $\mathsf{T}_{\gamma}$ be the fixed-point encoding defined in (BF). It holds that $\mathsf{T}_{\gamma}$ is $q$-Lipschitz where
$q=\sqrt{1-\gamma(2\alpha-\gamma L^{2}_{\alpha,\beta})},$
uniformly over $w\in\mathcal{H}$. In particular, $\mathsf{T}^{(w,\beta)}_{\gamma}$ is a contraction when $\gamma\in(0,2\alpha/L^{2}_{\alpha,\beta})$.
Proof.
The proof is similar to the one of Lemma 4.1. $\blacksquare$
Douglas-Rachford splitting.
Analogously, we can also consider the Douglas-Rachford splitting [lions_splitting_1979], whose operator $\mathsf{T}$ is given by
$\mathsf{T}^{(w,\beta)}=\frac{1}{2}(\operatorname{Id}+R_{\Phi^{w}_{\alpha,\beta}}\circ R_{\mathsf{A}}).$ (DR)
where $R_{\bullet}$ is the reflected resolvent given by $R_{\bullet}=2\operatorname{\mathsf{J}}_{\bullet}-\operatorname{Id}$. Note that $\operatorname{\mathsf{J}}_{\mathsf{A}}(\operatorname{Fix}(\mathsf{T}))=\operatorname{Zer}(\mathsf{A}+\Phi_{\alpha,\beta}^{w})$, such that $\mathsf{Z}=\operatorname{\mathsf{J}}_{\mathsf{A}}$ takes over the role of the fixed-point transportation map.
Lemma 4.3.
[moursi_douglas_2019, Corollary 3.1] Assume Assumption 7 holds and let $\mathsf{T}$ be the fixed-point encoding defined in (DR). It holds that $\mathsf{T}^{(w,\beta)}$ is $q$-Lipschitz where
$q=\frac{1}{2}+\frac{1}{2}\left(\frac{1-2\alpha+L_{\alpha,\beta}^{2}}{1+2\alpha+L_{\alpha,\beta}^{2}}\right)^{1/2},$
uniformly over $w\in\mathcal{H}$. In particular, $\mathsf{T}^{(w,\beta)}$ is a contraction.
4.2 Three-Operator Splitting Scheme
In this section we focus on problems of the form (P) with $\mathsf{M}=\mathsf{F}+\mathsf{A}+\mathsf{B}$, namely
$\operatorname{VI}(\mathsf{G},\mathcal{S}_{0}),\qquad\text{where }\mathcal{S}_{0}\triangleq\operatorname{Zer}(\mathsf{F}+\mathsf{A}+\mathsf{B}).$ (P-3-Split)
We refine the problem assumptions through the following assumption:
Assumption 8.
We assume the following on Problem (P-3-Split):
  1. 1.
    the set $\mathcal{S}_{0}$ is nonempty;
  2. 2.
    $\mathsf{G}:\mathcal{H}\to\mathcal{H}$ is monotone and $L_{\mathsf{G}}$-Lipschitz continuous;
  3. 3.
    $\mathsf{F}\colon\mathcal{H}\to\mathcal{H}$ is $L_{\mathsf{F}}$-Lipschitz continuous;
  4. 4.
    $\mathsf{A},\mathsf{B}\colon\mathcal{H}\to 2^{\mathcal{H}}$ are maximally monotone such that $\mathsf{A}$ has bounded domain satisfying $\operatorname{dom}(\mathsf{A})\subset\operatorname{dom}(\mathsf{B})$.
Under these assumptions, $\operatorname{dom}(\mathsf{M})=\operatorname{dom}(\mathsf{A})$. We note that in most practical examples, $\operatorname{dom}(\mathsf{B})=\mathcal{H}$, such that $\operatorname{dom}(\mathsf{A})\subset\operatorname{dom}(\mathsf{B})$ holds naturally.
We consider the auxiliary function $\Psi^{w}_{\alpha,\beta}:\mathcal{H}\to\mathcal{H}$ given by
$\Psi^{w}_{\alpha,\beta}(v):=\mathsf{F}(v)+\beta\mathsf{G}(v)+\alpha(v-w),$
which is $\alpha$-strongly monotone and $L_{\alpha,\beta}\triangleq L_{\mathsf{F}}+\beta L_{\mathsf{G}}+\alpha$-Lipschitz continuous, and hence $\nu_{\alpha,\beta}=\alpha/L_{\alpha,\beta}^{2}$-cocoercive. The auxiliary Problem (Aux) reads as $\operatorname{Zer}(\mathsf{A}+\mathsf{B}+\Psi^{w}_{\alpha,\beta})$, and may be solved through a three-operator splitting scheme [davis_threeoperator_2017a], namely through the fixed-point encoding map $\mathsf{T}^{(w,\beta)}_{\gamma}\colon\mathcal{H}\to\mathcal{H}$ depending on a step-size $\gamma$ given by
$\mathsf{T}^{(w,\beta)}_{\gamma}=\operatorname{Id}-\operatorname{\mathsf{J}}_{\gamma\mathsf{B}}+\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}\circ(2\operatorname{\mathsf{J}}_{\gamma\mathsf{B}}-\operatorname{Id}-\gamma\Psi^{w}_{\alpha,\beta}\circ\operatorname{\mathsf{J}}_{\gamma\mathsf{B}}).$ (TOS)
We know that $\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}(\operatorname{Fix}(\mathsf{T}))=\operatorname{Zer}(\mathsf{A}+\mathsf{B}+\Psi_{\alpha,\beta}^{w})$, such that $\mathsf{Z}=\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}$ is the fixed-point transportation map. The operator $\mathsf{T}^{(w,\beta)}_{\gamma}$ is non-expansive under the provided assumptions, and a contraction for well-chosen $\gamma$.
Lemma 4.4.
[davis_threeoperator_2017a, Theorem D.6] Let $\mathsf{T}^{(w,\beta)}_{\gamma}$ be the fixed-point encoding defined in (TOS). It holds that $\mathsf{T}^{(w,\beta)}_{\gamma}$ is non-expansive. Moreover, if $\mathsf{B}$ is $L_{\mathsf{B}}$-Lipschitz continuous and $\gamma<\eta\nu_{\alpha,\beta}$, then, uniformly over $w\in\mathcal{H}$, $\mathsf{T}^{(w,\beta)}_{\gamma}$ is $\sqrt{1-q}$-Lipschitz with constant
$q=\frac{2\gamma\alpha(1-\eta)}{(1+\gamma L_{B})^{2}},$
where $\eta\in[0,1]$ is arbitrary. In particular, $\mathsf{T}^{(w,\beta)}_{\gamma}$ is a contraction when
$\gamma\in\left(0,\frac{2\alpha(1-\eta)-2L_{\mathsf{B}}+\sqrt{4\alpha^{2}(1-\eta^{2})-8\alpha(1-\eta)L_{\mathsf{B}}}}{2L_{\mathsf{B}}^{2}}\right).$
Remark 4.2.
Although we require $\mathsf{B}$ to be Lipschitz continuous for Assumption 2 to be satisfied, we shall allow it not to be in our implementations, as per Remark 3.2. This is specifically the case in Section 5.3.
5 Numerical Experiments
In this section, we develop numerical examples showcasing the versatility and applicability of Algorithm 1. Specifically, we consider an equilibrium problem in Section 5.1, a least-norm least-squares problem in Section 5.2, and an image inpainting problem in Section 5.3.
All the code may be found on the author’s GitHub page [^1] See https://github.com/Hierarchical-VIs/Regularisation-Methods-for-HVIs., and is run on Intel Xeon Platinum 8380 CPUs.
5.1 Equilibrium selection
We consider the two-player zero-sum game from [samadi_improved_2025], given by
$\begin{dcases}&\min_{x_{1}}f_{1}(x_{1},x_{2})\triangleq 20-0.1x_{1}x_{2}+x_{1}\\ &\text{s.t.}~x_{1}\in X_{1}\triangleq[11,60]\end{dcases}\hskip 17.00024pt\begin{dcases}&\min_{x_{2}}f_{2}(x_{1},x_{2})\triangleq-20+0.1x_{1}x_{2}-x_{1}\\ &\text{s.t.}~x_{2}\in X_{2}\triangleq[10,50].\end{dcases}$
Specifically, we seek a saddle point $(x_{1}^{*},x_{2}^{*})\in X\triangleq X_{1}\times X_{2}$ of the function $f(x_{1},x_{2})\triangleq 20-0.1x_{1}x_{2}+x_{1}$, namely a point that satisfies
$f(x_{1}^{*},x_{2})\leq f(x_{1}^{*},x_{2}^{*})\leq f(x_{1},x_{2}^{*})\qquad\text{for all $(x_{1},x_{2})\in X$.}$
The solution set is characterized through the inclusion
$0\in\mathsf{F}(x_{1},x_{2})+\operatorname{\mathsf{NC}}_{X}(x_{1},x_{2})\triangleq\begin{bmatrix}\nabla_{x_{1}}f(x_{1},x_{2})\\ -\nabla_{x_{2}}f(x_{1},x_{2})\end{bmatrix}+\operatorname{\mathsf{NC}}_{X}(x_{1},x_{2})=\begin{bmatrix}0&-0.1\\ 0.1&0\end{bmatrix}\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}+\begin{bmatrix}1\\ 0\end{bmatrix}+\operatorname{\mathsf{NC}}_{X}(x_{1},x_{2})$
Analytically, we know the set of solutions to be $X_{1}\times\{10\}$. We attempt to find the best solution, namely to solve the problem
$\min\left\{\phi(x_{1},x_{2})\colon(x_{1},x_{2})\in\operatorname{Zer}(\mathsf{F}+\operatorname{\mathsf{NC}}_{X})\right\},$
where $\phi(x)=\tfrac{1}{2}\|x\|^{2}$, whose analytical solution is $(11,10)$. Equivalently, we aim to find a point $(x_{1}^{*},x_{2}^{*})\in\operatorname{Zer}(\mathsf{F}+NC_{X})$ such that
$\langle\nabla\phi(x_{1}^{*},x_{2}^{*}),(w_{1},w_{2})-(x_{1}^{*},x_{2}^{*})\rangle\geq 0\qquad\text{for all $(w_{1},w_{2})\in\operatorname{Zer}(F+\operatorname{\mathsf{NC}}_{X})$}.$
This matches Problem (P) with $\mathsf{G}=\nabla\phi$ for the upper level operator, and $\mathsf{M}=\mathsf{F}+\operatorname{\mathsf{NC}}_{X}$ for the lower level operator.
Experimental Setup.
For all the experiments, we consider the relaxation parameter of the inner loop $\theta_{k}\equiv\theta=0.7$ to be constant. Across inner loops, we consider the acceleration parameter $\tau_{k}\equiv\tau$ to be constant and to be the largest value satisfying Equation (3.1). We assume $\beta_{n}=(n+1)^{-\eta}$, where $\eta=0.55$, and that $\varepsilon_{n}=\overline{\varepsilon}\cdot{(n+1)^{-2}}$, where $\overline{\varepsilon}=10^{-3}$. We run a total of $1000$ iterations.
Results.
Figure 2 shows the averaged iterates $(\overline{w}_{n})$, using Forward-Backward for the auxiliary problem, for various starting points, along with the feasible region $X$. Recall that, although we call it the feasible region, we have no a priori guarantee that the iterates remain within said region.
Figure 2: Average iterates for various initial points and proximal parameters.
Moreover, we note that the optimality and feasibility gaps may be computed through a small subproblem in this simple example. Figures 3 and 4 plot these values, for the different initial points. Although we observe differences, we note that all curves are decreasing to 0. We note the importance of the tuning of the proximal parameter $\alpha$. In the present example, a smaller value yields faster convergence. As will be evidenced in later simulations, this is not always the case.
Figure 3: Optimality gap $\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{n})$ for various initial points and proximal parameters.
Figure 4: Feasibility gap $\operatorname{Gap}_{\operatorname{feas}}(\overline{w}_{n})$ for various initial points and proximal parameters.
In Figures 5 and 6, we compare various fixed-point encodings of the auxiliary problem. Specifically, we implement the methods from Section 4.1, namely forward-backward (FB), backward-forward (BF) and Douglas-Rachford (DR). We set $\overline{\varepsilon}=1$ and $\alpha=0.1$. We observe different behaviors for the different algorithms, but do not focus further on these in this work.
Figure 5: Optimality gap $\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{n})$ for various auxiliary problem fixed-point encodings.
Figure 6: Feasibility gap $\operatorname{Gap}_{\operatorname{feas}}(\overline{w}_{n})$ for various auxiliary problem fixed-point encodings.
5.2 Least-Norm Least-Squares
Inspired by [staudigl2023random], for a matrix $A\in\mathbb{R}^{P\times Q}$ and a vector $b\in\mathbb{R}^{P}$, we consider the bilevel optimization problem given by
$\min_{u\in\mathbb{R}^{Q}}\left\{\tfrac{1}{2}\|u\|^{2}~\colon~u\in\operatorname*{argmin}_{v\in X}\left\{\tfrac{1}{2}\|{Av-b}\|^{2}\right\}\right\},$
where $X=[-1000,1000]^{Q}$. We can reformulate this as a hierarchical variational inequality problem of the form (P-2-Split) with $\mathsf{G}(u)=u$, $\mathsf{A}=\partial\delta_{X}=\operatorname{\mathsf{NC}}_{X}$, and $\mathsf{F}(v)=2A^{\top}(Av-b)$. We will make use of the Forward-Backward operator for the inner loop iterations.
Problem Setup.
We test the algorithm on four randomly generated instances with different dimensions given by $(P,Q)\in\{(70,100),(100,200),(100,500),(300,500)\}$. For each instance, the matrix $A$ is generated as a low-rank (and sparse) matrix via $A=U_{1}\cdot U_{2}$, where $U_{1}\in\mathbb{R}^{P\times R}$ and $U_{2}\in\mathbb{R}^{R\times Q}$ are chosen such that each component follows a standard normal law, with $R=50$ to induce a low-rank structure. Moreover, to ensure the eigenvalues are not too large such that the step-size is not too small, we clip each singular value to the interval $[0,10]$. A vector $s\in\mathbb{R}^{Q}$ is randomly generated with $20$ non-zero entries according to a uniform law on $[0,10]$, and the observation vector $b$ is constructed as $b=As+\nu$, where $\nu$ is a normal random vector with small entries, namely a standard normal vector scaled by a factor $0.1$. During the random generation we ensure that $z=A^{\dagger}b\in X$, such that the analytical solution is known to be $z$, where $A^{\dagger}$ is the Moore-Penrose inverse.
Experimental Setup.
For the inner loop, we set $\theta_{k}\equiv\theta=0.75$ to be constant. Across inner loops, we consider $\tau_{k}\equiv\tau$ to be constant and to be the largest value satisfying Equation (3.1). We set $\beta_{n}=(n+1)^{-\eta}$, where $\eta=0.55$, and $\varepsilon_{n}=\overline{\varepsilon}\cdot(n+1)^{-1}$, where $\overline{\varepsilon}=\alpha\cdot 10^{-3}$. We run a total of $2000$ iterations, with an initial point $w_{0}$ randomly generated according to a standard normal distribution scaled by a factor $0.1$.
Results.
Figures 7 and 8 show the evolution of the error sequence $\|w_{n}-z\|$ and the lower-level function gap sequence $\tfrac{1}{2}\|Aw_{n}-b\|^{2}-\tfrac{1}{2}\|Az-b\|^{2}$ across iterations. As expected, we observe a decrease for each problem. We note that the initial decrease for the error sequence is consequential and stagnates for a number of iterations afterwards. In fact, for $\alpha=0.1$, the prescribed $2000$ iterations are not enough to observe a second decrease phase, whereas it is sufficient for $\alpha\in\{1,10\}$.
Figure 7: Evolution of error sequence $\|w_{n}-z\|$ for various problem dimensions and various proximal parameters.
Figure 8: Evolution of lower-level function gap sequence $\tfrac{1}{2}\|Aw_{n}-b\|^{2}-\tfrac{1}{2}\|Az-b\|^{2}$ for various problem dimensions and various proximal parameters.
5.3 Image Inpainting
Finally, we consider the image inpainting problem, as done in [maulen_inertial_2024, Cortild:2024aa]. We represent a grayscale image $Y$ of dimension $(P,Q)$ by a matrix in $X\triangleq[0,1]^{P\times Q}\subset\mathcal{H}\triangleq\mathbb{R}^{P\times Q}$. We denote by $\Omega\in\{0,1\}^{P\times Q}$ a mask such that $\Omega_{i,j}=0$ indicates that the pixel at position $(i,j)$ has been damaged. We denote by $\mathcal{R}$ the linear operator that maps an image to an image whose elements in $\Omega$ have been corrupted, namely $\mathcal{R}\colon\mathcal{H}\to\mathcal{H},Y\mapsto\tilde{Y}$ with $\tilde{Y}_{i,j}=\Omega_{i,j}\cdot Y_{i,j}$. We note that $\mathcal{R}$ is a self-adjoint bounded linear operator with operator norm $1$. We define the corrupt image by $Y_{\text{corrupt}}\triangleq\mathcal{R}(Y)$. The aim of the image inpainting problem is to recover the original image $Y$ from the corrupt image $Y_{\text{corrupt}}$ and the corruption map $\mathcal{R}$. Mathematically, we consider the problem
$\min_{Y\in X}\left\{\tfrac{1}{2}\|\mathcal{R}(Y)-Y_{\text{corrupt}}\|^{2}+\sigma\|Y\|_{*}\right\},$
where $\|\cdot\|_{*}$ denotes the nuclear norm, which prones smoothness within the image, and $\sigma$ is a regularization parameter. Specifically, we seek the least-norm solution to the above problem, given by
$\min_{Y\in\mathcal{H}}\left\{\tfrac{1}{2}\|Y\|^{2}~\colon~Y\in\operatorname*{argmin}_{Y\in X}\left\{\tfrac{1}{2}\|\mathcal{R}(Y)-Y_{\text{corrupt}}\|^{2}+\sigma\|Y\|_{*}\right\}\right\}.$
This matches Problem (P-3-Split) with $\mathsf{G}=\operatorname{Id}$, $\mathsf{A}=\sigma\partial\|\cdot\|_{*}$, $\mathsf{B}=\operatorname{\mathsf{NC}}_{X}$, and $\mathsf{F}(Y)=\mathcal{R}^{*}(\mathcal{R}(Y)-Y_{\text{corrupt}})$. We note that Assumption 8 is satisfied, and that the operator $\mathsf{T}$ is nonexpansive per Lemma 4.4. Although this setting does not strictly conform to our theoretical findings, its convergence properties are expected to be analogous due to the finite-dimensional nature of the problem, as detailed in Remark 4.2.
Experimental Setup.
For the inner loop, we set $\theta_{k}\equiv\theta=0.75$ and we consider $\tau_{k}\equiv\tau$ to be constantly the largest value satisfying Equation (3.1) across every outer iteration. We set $\beta_{n}=(n+1)^{-\eta}$ for $\eta=0.55$ and $\varepsilon_{n}=\overline{\varepsilon}\cdot(n+1)^{-2}$ for $\overline{\varepsilon}=2$. We run a total of $5000$ iterations, with an initial point $w_{0}=Y_{\text{corrupt}}$. We set the regularization parameter to be $\sigma=50$, and generate $\Omega$ randomly such that $20\%$ of the image is corrupted.
Results.
Figure 9 shows the original image, the corrupt image and the image restored using the above procedure for various proximal parameters. Table 1 reports the total number of inner iterations run, along with the time spent to produce the results. Figure 10 plots the lower-level objective across outer iterations for each value of $\alpha$, which is expected to converge to the lower-level objective value and not to 0, as seen in the plot. We observe a loss in the solution quality for $\alpha=100$, although it is the fastest method. We also note that the solution qualities are very similar for $\alpha=0.1$ and $\alpha=1$, but the latter requires a larger number of total iterations.
Figure 9: Result of image inpainting procedure for various proximal parameters.
Table 1: Number of inner iterations and clocktime to produce Figure 9.
$\alpha=0.1$ $\alpha=1$ $\alpha=10$ $\alpha=100$
Total inner iterations 495096 156365 92867 73901
Time in minutes 294 80 38 36
Figure 10: Evolution of lower-level objective function $\tfrac{1}{2}\|\mathcal{R}(Y)-Y_{\text{corrupt}}\|^{2}+\sigma\|Y\|_{*}$ for various proximal parameters.
6 Conclusion
In this work, we proposed a double loop path following method for solving a hierarchical variational inequalities. The inner loop employs a Krasnoselskii-Mann type iteration for solving a suitable auxiliary problem. The outer loop takes the form of a restarting procedure, giving rise to a diagonal equilibrium tracking method (DANTE). The obtained convergence rates are comparable with the state of the art results, obtained for specific numerical schemes. Our analysis is very flexible with respect to the chosen fixed-point encoding strategy of the lower level solution. Importantly, our rates are independent of the specific fixed-point encoding, as long as the stated assumptions are satisfied. We belief that our work opens up the possibilities for various interesting directions for future research. First, it would be very interesting to device a fully adaptive scheme without restarts. Second, an important future extension is the inclusion of stochastic data. All these directions are the subject of ongoing research.
Acknowledgments. Daniel Cortild acknowledges the support of the Clarendon Funds Scholarships. Meggie Marschner and Mathias Staudigl acknowledge financial support from the PGMO and the Deutsche Forschungsgemeinschaft (DFG) - Projektnummer 556222748 "non-stationary hierarchical minimization".

### v2

@@
 This class of problems is an extremely active area of research recently, driven by important practical applications in engineering, economics and machine learning.
-For recent comprehensive reviews, see [dempe2020bilevel, liu2021investigating].
+Comprehensive reviews can be found in [dempe2020bilevel, liu2021investigating].
 A particularly appealing class of bi-level optimization problems are those in which only a single decision variable is involved at both layers of the problem.
@@
 $\operatorname{\mathsf{NC}}_{\mathcal{K}}(v)\triangleq\left\{\begin{array}[]{ll}\varnothing&\text{if }v\notin\mathcal{K}\\ \{p\in\mathcal{H}|\langle p,u-v\rangle\leq 0,\;\forall u\in\mathcal{K}\}&\text{ else.}\end{array}\right$.
-Hence, if we set the operator $\mathsf{M}=\mathsf{F}+\operatorname{\mathsf{NC}}_{\mathcal{K}}$, Problem (P) specializes to a hierarchical variational inequality problem given by
+Hence, by setting $\mathsf{M}=\mathsf{F}+\operatorname{\mathsf{NC}}_{\mathcal{K}}$, Problem (P) specializes to a hierarchical variational inequality problem
 $\operatorname{VI}(\mathsf{G},\mathcal{S}_{0}),\qquad\text{where $\mathcal{S}_{0}$ is the solution set of $\operatorname{VI}(\mathsf{F},\mathcal{K})$}.$
-For recent studies of such problems in finite-dimensional euclidean spaces, we refer to [alves_inertial_2025] and [samadi_improved_2025].
+For recent studies of such problem formulations in finite-dimensional Euclidean spaces, we refer to [alves_inertial_2025] and [samadi_improved_2025].
 Example 1.2 (Equilibrium Selection).
@@
 If $g$ is convex and Fréchet differentiable, this becomes a hierarchical variational inequality with $\mathsf{G}=\nabla g$, hence a particular instance of Problem (P).
-This viewpoint covers a broad range of models encountered in, e.g., engineering, economics, inverse problems and control (see [hintermuller2011first, hintermuller2015bi-level, hintermuller2014several, outrata2000generalized, de2023bi-level]).
+This viewpoint covers a broad range of models encountered in, e.g., engineering, economics, inverse problems, and control (see [hintermuller2011first, hintermuller2015bilevel, hintermuller2014several, outrata2000generalized, de2023bilevel]).
 1.1 Contributions
 Numerical schemes with complexity statements for resolving Problem (P) have been recently published in the papers [Kaushik:2021aa, samadi_improved_2025] and [alves_inertial_2025].
-[Kaushik:2021aa] extend the simple bi-level setting to a lower level problem defined in terms of a variational inequality while keeping the upper level problem as a convex optimization problem.
-This is a special instance of Problem (P), which is related to mathematical programming with equilibrium constraints [Luo_Pang_Ralph_1996], and to the equilibrium selection problem in game theory [Benenati:2022aa, Benenati:2023aa].
-The recent papers [samadi_improved_2025, alves_inertial_2025] go one step further and study the special instance of Problem (P) in which the upper and lower level problem are defined in terms of a variational inequality problem.
-In this paper we generalize the aforementioned existing results by considering a more general class of hierarchical equilibrium problems admitting the formulation (P).
-We prove rates on both levels of the equilibrium problem via a very flexible numerical scheme, which can easily be modulated with respect to problem structure (specifically of the lower-level inclusion problem).
+[Kaushik:2021aa] extend the simple bi-level setting to a lower level problem defined in terms of a variational inequality, while keeping the upper level problem as a convex optimization problem.
+This is a particular case of Problem (P), which is related to mathematical programming with equilibrium constraints [Luo_Pang_Ralph_1996], and to the equilibrium selection problem in game theory [Benenati:2022aa, Benenati:2023aa].
+The recent papers [samadi_improved_2025, alves_inertial_2025] go one step further and study a special instance of Problem (P) in which the upper- and lower-level problems are defined in terms of a variational inequality.
+In this paper we generalize the aforementioned existing results, by considering a significantly larger class of hierarchical equilibrium problems, admitting the formulation (P).
+We prove rates on both levels of the equilibrium problem via a very flexible numerical scheme, which can easily be modulated to prevailing problem structure (specifically of the lower-level inclusion problem).
 Our main contributions are threefold:
@@
   • Monotone-inclusion lower level and modular oracle.
-Rather than considering a variational inequality at the lower level, we allow the feasible set to be defined as the solution set of an abstract monotone inclusion.
-This formulation allows to draw from the rich theory of monotone operators and enables a very general algorithmic design based on a monotone-inclusion oracle obtained through a fixed-point encoding, which can be tailored to the structure of the operators involved.
+Rather than considering a variational inequality at the lower level, we allow the feasible set to be defined as the solution set of an abstract monotone inclusion problem.
+Within this general framework, we adopt a very general algorithmic design based on a monotone-inclusion oracle obtained through a fixed-point encoding, which can be tailored to the structure of the operators involved.
+As a result, our numerical scheme and our analysis is applicable to entire families of relevant splitting methods, and not just a single one.
   • Quantitative guarantees via gap functions.
 While much of the existing literature focuses on weak convergence results without explicit rate statements, we introduce gap functions for both the lower- and upper-level problems and establish quantitative convergence guarantees for these measures.
-Our results are formulated within a specific geometric setting in which we assume that the feasibility gap function is assumed to satisfy an Hölderian error bound condition such as [Cabot2005] in the context of hierarchical optimization and more recently [samadi_improved_2025] and [boct2025accelerating].
-We achieve these results via the careful design of an algorithmic template drawing inspiration from classical diagonal methods involving iterative Tikhonov and proximal regularization via anchor terms.
+Our results are formulated within a specific geometric setting in which we assume that the gap function associated with the lower-level problem satisfies a Hölderian error bound condition.
+Such settings have already been considered in [Cabot2005] in the context of hierarchical optimization, and more recently in [samadi_improved_2025] and [boct2025accelerating].
+We achieve these results via the careful design of an algorithmic template drawing inspiration from classical diagonal methods, involving iterative Tikhonov and proximal regularization via anchor terms.
 By constructing a family of auxiliary combined equilibrium problems, our method constructs a sequence in an inner loop method, which is designed to track a sequence of temporal solutions.
-This procedure is essentially an inertial Krasnoselskii-Mann iteration and thus only relies on the availability of a ready-to-compute fixed-point encoding map for the lower-level solution set $\mathcal{S}_{0}=\operatorname{Zer}(\mathsf{M})$.
-An outer-loop procedure updates the Tikhonov parameter and the anchor term to steer the trajectory of the numerical algorithm towards a solution of the entire Problem (P).
-Diagonal schemes for solving convex programs are a classical topic [alart1991penalization].
+This procedure is essentially an inertial Krasnoselskii-Mann (KM) iteration and thus only relies on the availability of a ready-to-compute fixed-point encoding map for the lower-level solution set $\mathcal{S}_{0}=\operatorname{Zer}(\mathsf{M})$.
+An outer-loop restarting procedure updates the Tikhonov parameter and the anchor term to steer the trajectory of the numerical algorithm towards a solution of (P).
+Diagonal schemes for solving convex programs are rather classical (see e.g.
+[alart1991penalization]).
 The idea of extending this approach to hierarchical equilibrium problems is taken from [Facchinei:2014aa] and [Lampariello:2020aa].
-In particular, [Lampariello:2020aa] also proposes a kind of double-loop scheme with a forward-backward iteration acting in the inner loop.
+In particular, [Lampariello:2020aa] also proposes a kind of double-loop scheme with a forward-backward algorithm acting in the inner loop.
 In this paper, we consider a more general class of splitting problems, with explicit rate guarantees at both levels of the problem.
-Specifically, in the case where $\mathsf{G}$ is merely monotone (opposed to strongly monotone), we obtain weak convergence of the averaged sequence of iterates $(\overline{w}_{N})$, with rates of the form, for arbitrary $b\in(0,1)$,
+Specifically, in the case where $\mathsf{G}$ is merely monotone (as opposed to strongly monotone), we obtain weak convergence of the averaged sequence of iterates $(\overline{w}_{n})_{n\in\mathbb{N}}$, with rates of the form, for arbitrary $b\in(0,1)$,
 $\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{n})\leq\tilde{\mathcal{O}}\left(\frac{1}{n^{1-b}}\right),\qquad\text{and}\qquad 0\leq\operatorname{Gap}_{\operatorname{feas}}(\overline{w}_{n})\leq\tilde{\mathcal{O}}\left(\frac{1}{n^{\min(b,1-b)}}\right),$
 for suitably defined optimality and feasibility gaps (see Section 3.2 for precise definition).
-In this setting, $\tilde{\mathcal{O}}$ disregards logarithmic factors, which are natural consequences of the double-loop structure of our algorithm.
-Ignoring these logarithmic factors, which previous works do not obtain due to single-loop structures, our work is on par with established benchmarks:
+In this setting, the integer $n$ measures the number of restarts, and $\tilde{\mathcal{O}}$ disregards logarithmic factors, which are consequences of the double-loop structure of our algorithm.
+Besides these logarithmic factors, our work is on par with published single-loop algorithms:
   • In [Kaushik:2021aa, Corollary 3.5], the authors study optimization problems with Cartesian variational inequality constraints, where the objective is convex and the constraint set is a Cartesian product associated with a monotone mapping.
 Their algorithm resembles a forward-backward method, cast into a single-loop scheme.
-For the deterministic variant of their method, they obtain rates of $\mathcal{O}(1/N^{0.5-b})$ for the optimality gap and $\mathcal{O}(1/N^{b})$ for the feasibility gap, for $b\in(0,0.5)$.
-Therefore, ignoring the logarithmic factors, the rates achieved in the present work are more favorable.
+For the deterministic variant of their method, they obtain rates of $\mathcal{O}(1/N^{0.5-b})$ for the optimality gap and $\mathcal{O}(1/N^{b})$ for the feasibility gap, for $b\in(0,0.5)$, $N$ being the iteration count.
+Therefore, ignoring logarithmic factors, the rates achieved in the our work are more favorable.
   • In [samadi_improved_2025, Theorem 3.6], the authors study a regularized extragradient method, and establish rates of $\mathcal{O}(1/N^{1-b})$ for the optimality gap and $\mathcal{O}(1/N^{b})$ for the feasibility gap.
@@
 When $\mathsf{G}$ is strongly monotone, our rates improve to
-$\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{N})\leq\tilde{\mathcal{O}}\left(\frac{1}{N}\right),\qquad\text{and}\qquad 0\leq\operatorname{Gap}_{\operatorname{feas}}(\overline{w}_{N})\leq\tilde{\mathcal{O}}\left(\frac{1}{N}\right).$
-These rates are comparable to those in [samadi_improved_2025, Theorem 4.6], where a regularized extragradient method is analyzed.
-To the best of our knowledge, no other works address hierarchical programming problems in the strongly monotone setting in a manner that allows a direct comparison.
+$\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{N})\leq\tilde{\mathcal{O}}\left(\frac{1}{N}\right),\qquad\text{and}\qquad 0\leq\operatorname{Gap}_{\operatorname{feas}}(\overline{w}_{N})\leq\tilde{\mathcal{O}}\left(\frac{1}{N}\right),$
+where, again, $N$ measures the number of restarts. These rates are comparable to those in [samadi_improved_2025, Theorem 4.6], where a regularized extragradient method is analyzed.
 Organization of the paper.
@@
 1.2 Notation and Definitions
+We adopt standard notation and terminology from variational analysis, as in [BauCom16].
 Let $\mathcal{H}$ be a real Hilbert space with inner product $\langle\cdot,\cdot\rangle$ and corresponding norm $\|{\cdot}\|$.
-A set-valued operator $\mathsf{A}\colon\mathcal{H}\to 2^{\mathcal{H}}$ is monotone if $\langle u-v,b-a\rangle\geq 0$ for all $(u,b),(v,a)\in\operatorname{Gr}(\mathsf{A})\triangleq\{(v,a)\in\mathcal{H}\times\mathcal{H}|a\in\mathsf{A}(v)\}$.
+A set-valued operator $\mathsf{A}\colon\mathcal{H}\to 2^{\mathcal{H}}$ is $\mu$-monotone if $\mu\geq 0$ and
+$\langle b-a,u-v\rangle\geq\mu\|{u-v}\|^{2}\text{ for all }(u,b),(v,a)\in\operatorname{Gr}(\mathsf{A})\triangleq\{(v,a)\in\mathcal{H}\times\mathcal{H}|a\in\mathsf{A}(v)\}.$
+If $\mu=0$, we will say $\mathsf{A}$ is merely monotone.
 The operator $\mathsf{A}\colon\mathcal{H}\to 2^{\mathcal{H}}$ is maximally monotone if it is monotone and there exists no other monotone operator whose graph properly contains $\operatorname{Gr}(\mathsf{A})$.
-The resolvent of an operator $\mathsf{A}$ is defined as $\operatorname{\mathsf{J}}_{\mathsf{A}}=(\operatorname{Id}+\mathsf{A})^{-1}$.
-If $\mathsf{A}$ is maximally monotone, then the resolvent is a nonexpansive operator and thus single-valued.
-A mapping $\mathsf{T}\colon\mathcal{H}\to\mathcal{H}$ is quasi-contractive if $\|Tx-p\|\leq q\|x-p\|$ for all $p\in\text{Fix}(T)$ and some $q\in(0,1)$.
-We refer to [BauCom16] for more details on operator theory.
-For two sequences $(a_{n})\subset\mathbb{R}$ and $(b_{n})\subset\mathbb{R}$, we say that $a_{n}=O(b_{n})$ if there exists a constant $C>0$ such that $a_{n}\leq Cb_{n}$ for $n$ sufficiently large.
+The domain of an operator $\mathsf{A}$ is $\operatorname{dom}(\mathsf{A})=\{x\in\mathcal{H}|\mathsf{A}(x)\neq\varnothing\}$.
+A mapping $\mathsf{T}\colon\mathcal{H}\to\mathcal{H}$ is quasi-contractive if $\|Tx-p\|\leq q\|x-p\|$ for all $p\in\textnormal{Fix}(T)$ and some $q\in(0,1)$, where $\textnormal{Fix}(\mathsf{T})\triangleq\{x\in\mathcal{H}|x=\mathsf{T}(x)\}$ is the set of fixed points of $\mathsf{T}$.
+If the same inequality holds for $q=1$, then $\mathsf{T}$ is non-expansive.
+The resolvent of an operator $\mathsf{A}$ is defined as $\operatorname{\mathsf{J}}_{\mathsf{A}}\triangleq(\operatorname{Id}+\mathsf{A})^{-1}$.
+By Minty’s theorem, if $\mathsf{A}$ is maximally monotone, then the resolvent is a nonexpansive operator and thus single-valued.
+The normal cone mapping to a closed convex and non-empty set $\mathcal{K}\subset\mathcal{H}$ is defined as $\operatorname{\mathsf{NC}}_{\mathcal{K}}(x)=\{u\in\mathcal{H}|\langle u,y-x\rangle\leq 0\;\forall y\in\mathcal{K}\}$ for $x\in\mathcal{K}$ and $\operatorname{\mathsf{NC}}_{\mathcal{K}}(x)=\varnothing$ otherwise.
+Under convexity, $\operatorname{\mathsf{NC}}_{\mathcal{K}}$ is maximally monotone, and its resolvent coincides with the orthogonal projector $\Pi_{\mathcal{K}}(x)=\operatorname*{argmin}\{\|{y-x}\|:y\in\mathcal{K}\}$.
+For two sequences $(a_{n})\subset\mathbb{R}$ and $(b_{n})\subset\mathbb{R}$, we say that $a_{n}={\mathcal{O}}(b_{n})$ if there exists a constant $C>0$ such that $a_{n}\leq Cb_{n}$ for $n$ sufficiently large.
 We say that $a_{n}=o(b_{n})$ if $\lim_{n\to\infty}{b_{n}}/{a_{n}}=0$.
@@
 Assumption 1 guarantees that the overall Problem (P) has a solution. In particular, $\mathcal{S}_{0}$ is a convex and compact subset of $\mathcal{H}$ [BauCom16].
-Remark 2.2.
-We emphasize that our assumption includes the case $\mu=0$, in which $\mathsf{G}$ is only Lipschitz and monotone.
 For given parameters $(\alpha,\beta)\in(0,\infty)^{2}$ and an anchor point $w\in\mathcal{H}$, we define the associated auxiliary problem as finding
@@
 Our numerical scheme consists of two main procedures.
-The inner procedure is an inertial Krasnoselskii-Mann (KM) iteration, whose last iterate is designed to lie in a sufficiently small ball around the temporal solution $\overline{u}_{\beta}(w)$, whose radius is part of the parameters of the procedure.
-We use this last iterate as a new anchor point, and restart the KM iteration with fresh parameters.
-To this end, we assume access to a family of parameterized fixed-point mappings $\mathsf{T}_{k}\equiv\mathsf{T}^{(w,\beta)}_{k}\colon\mathcal{H}\to\mathcal{H}$, which satisfy the following condition.
+The inner procedure is an inertial KM iteration, whose last iterate is designed to lie in a sufficiently small ball around the temporal solution $\overline{u}_{\beta}(w)$, with iteratively defined radius.
+We subsequently use this last iterate as a new anchor point, and restart the KM iteration with fresh parameters.
+To this end, we assume to have access to a family of parameterized fixed-point mappings $\mathsf{T}_{k}\equiv\mathsf{T}^{(w,\beta)}_{k}\colon\mathcal{H}\to\mathcal{H}$, satisfying the following condition.
 Assumption 2.
@@
 Moreover, $\mathsf{T}_{k}^{(w,\beta)}$ is related to Problem (Aux) via a nonexpansive fixed-point transportation map $\mathsf{Z}^{(w,\beta)}\colon\mathcal{H}\to\mathcal{H}$ such that $\overline{u}_{\beta}(w)=\mathsf{Z}^{(w,\beta)}(p(w,\beta))$.
-Concrete examples of such operators are provided in Section 4. Generally, the operators may vary over the inner loop as a result of a varying step-size in the selected method.
+Concrete examples of such fixed-point encoding mappings are provided in Section 4.
 Remark 2.2.
 The proximal regularization parameter $\alpha>0$ is treated in our approach as an external input to the algorithm.
-Hence, its concrete numerical value used in the computation affects the central funnel $\overline{u}_{\beta}(w)$, as well as the fixed-point encoding map $\mathsf{T}^{(w,\beta)}_{k}$ and eventually also the fixed-point transportation map $\mathsf{Z}^{(w,\beta)}$.
+Hence, its concrete numerical value used in the computation affects the position of the solution $\overline{u}_{\beta}(w)$, the shape of the fixed-point encoding map $\mathsf{T}^{(w,\beta)}_{k}$, and eventually also the fixed-point transportation map $\mathsf{Z}^{(w,\beta)}$.
 However, since it is a fixed parameter, we simplify the notation and omit an explicit dependence from the involved mappings and operators on the value of $\alpha$.
 2.1 Inner Loop
-The inner loop of our tracking algorithm employs the procedure $\mathtt{KM}(v,\mathsf{T}^{(w,\beta)}_{\bullet},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$, see Algorithm 1, which is an inertial Krasnoselskii-Mann iteration with user-provided parameters embodied in terms of the initial point $v\in\mathcal{H}$, fixed-point encoding map $\mathsf{T}^{(w,\beta)}_{\bullet}$ satisfying Assumption 2, as well as momentum parameters $(\tau_{k})_{k}$ and relaxation parameters $(\theta_{k})_{k}$.
+The inner loop of our tracking algorithm employs the procedure $\mathtt{KM}(v,(\mathsf{T}^{(w,\beta)}_{k})_{k},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$, see Algorithm 1, which is an inertial Krasnoselskii-Mann iteration with user-provided parameters embodied in terms of the initial point $v\in\mathcal{H}$, fixed-point encoding mappings $(\mathsf{T}^{(w,\beta)}_{k})_{k}$ satisfying Assumption 2, as well as momentum parameters $(\tau_{k})_{k}$ and relaxation parameters $(\theta_{k})_{k}$.
 This fixed-point iteration is designed to return an approximation of the central funnel fixed-point $\overline{u}_{\beta}(w)$ via the fixed-point transportation map $\mathsf{Z}^{(w,\beta)}$.
-Algorithm 1 Function $\mathtt{KM}(v,\mathsf{T}^{(w,\beta)}_{\bullet},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$
-$v_{1}=v_{0}:=v$
+Algorithm 1 Function $\mathtt{KM}(v,(\mathsf{T}^{(w,\beta)}_{k})_{k},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$
+$v_{1}=v_{0}\triangleq v$
 Set $k=1$
@@
 Algorithm 1 employs the stopping time
-$\mathbb{K}(\varepsilon)\triangleq\inf\{k\geq 1|\;\|{v_{k+1}-z_{k}}\|\leq\varepsilon\}.$ (Stop)
-This stopping criterion is motivated by the following approximation result.
+$\mathbb{K}(\varepsilon)\triangleq\inf\{k\geq 1|\;\|{v_{k+1}-z_{k}}\|\leq\varepsilon\}.$ (2.1)
+This stopping criterion is motivated by the following approximation result, formulated in terms of general fixed-point encodings obeying the imposed assumptions.
 Lemma 2.1.
@@
 Denote by $p$ the common fixed-point of $(\mathsf{T}_{k})_{k}$.
-If $\mathbb{K}(\varepsilon)$ is finite and $v_{\mathbb{K}(\varepsilon)+1}$ is the result of Algorithm 1, then
+If the stopping time $\mathbb{K}(\varepsilon)$ defined in (2.1) is finite and $v_{\mathbb{K}(\varepsilon)+1}$ is the last iterate produced by Algorithm 1, then
 $\|{v_{\mathbb{K}(\varepsilon)+1}-p}\|\leq\mathtt{e}\triangleq\frac{\varepsilon}{\overline{\theta}}\left({1-\overline{\theta}}+\frac{\overline{q}}{1-\overline{q}}\right).$ (2.2)
 Proof.
-Given $\varepsilon>0$, denote $\mathbb{K}=\mathbb{K}(\varepsilon)$. By definition of $v_{k+1}$ and quasi-contractiveness of $\mathsf{T}_{k}$, we get
+Given $\varepsilon>0$, let $\mathbb{K}=\mathbb{K}(\varepsilon)$. By definition of $v_{k+1}$ and quasi-contractiveness of $\mathsf{T}_{k}$, we get
 $\|{z_{k}-p}\|$ $=\|{z_{k}-\mathsf{T}_{k}(z_{k})+\mathsf{T}_{k}(z_{k})-p}\|\leq\|{\frac{v_{k+1}-z_{k}}{\theta_{k}}}\|+\|{\mathsf{T}_{k}(z_{k})-p}\|$
@@
 $\leq\varepsilon\frac{1-\theta_{\mathbb{K}}}{\theta_{\mathbb{K}}}+\overline{q}\cdot\|{z_{\mathbb{K}}-p}\|$
-$\leq\frac{\varepsilon}{\theta_{\mathbb{K}}}\left({1-\theta_{\mathbb{K}}}+\frac{\overline{q}}{1-\overline{q}}\right)\leq\frac{\varepsilon}{\overline{\theta}}\left({1-\overline{\theta}}+\frac{\overline{q}}{1-\overline{q}}\right),$
-as wanted. $\blacksquare$
+$\leq\frac{\varepsilon}{\theta_{\mathbb{K}}}\left({1-\theta_{\mathbb{K}}}+\frac{\overline{q}}{1-\overline{q}}\right)\leq\frac{\varepsilon}{\overline{\theta}}\left({1-\overline{\theta}}+\frac{\overline{q}}{1-\overline{q}}\right).$
+$\blacksquare$
 Lemma 2.1 has important consequences for the tracking properties of the anchor points $w$ which we recursively generate with the restarting procedure explained next.
-To give an outlook, consider the sequence $(v_{k})_{k}$ produced by the function $\mathtt{KM}(v,\mathsf{T}^{(w,\beta)}_{\bullet},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$, and let $w^{+}\triangleq\mathsf{Z}^{(w,\beta)}(v_{\mathbb{K}(\varepsilon)+1})$.
+To give an outlook, consider the sequence $(v_{k})_{k}$ produced by the function $\mathtt{KM}(v,(\mathsf{T}^{(w,\beta)}_{k})_{k},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$, and let $w^{+}\triangleq\mathsf{Z}^{(w,\beta)}(v_{\mathbb{K}(\varepsilon)+1})$.
 Lemma 2.1 guarantees that this point satisfies the tracking guarantee
 $\|{w^{+}-\overline{u}_{\beta}(w)}\|\leq\mathtt{e}=\frac{\varepsilon}{\overline{\theta}}\left({1-\overline{\theta}}+\frac{\overline{q}}{1-\overline{q}}\right),$ (2.3)
-by nonexpansiveness of $\mathsf{Z}^{(w,\beta)}$. The upper bound is thus of order $O(\varepsilon)$, and thus controlled by the user-defined targeted precision.
+by nonexpansiveness of $\mathsf{Z}^{(w,\beta)}$. The upper bound is of order $\mathcal{O}(\varepsilon)$, and thus controlled by the user-defined targeted precision.
 2.2 Restarting Procedure
-Given an anchor point $w$, Algorithm 1 produces a sequence of points $(v_{k})$ whose last iterate corresponds to the output of the function $\mathtt{KM}(v,\mathsf{T}^{(w,\beta)}_{\bullet},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$ and which provides a good candidate for a new anchor point, given the tracking property (2.2).
-Repeating this over time for a fixed number of iterations is the key updating step constituting our outer loop procedure.
-The mechanics of this procedure gives rise to an diagonal equilibrium tracking method (DANTE), formally described in Algorithm 2.
+Given an anchor point $w$, Algorithm 1 produces a sequence of points $(v_{k})_{k}$, whose last iterate corresponds to the output of the function $\mathtt{KM}(v,(\mathsf{T}^{(w,\beta)}_{k})_{k},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$.
+In view of the tracking property established in (2.3), we will use this point as the new anchor point for the inner-loop KM iteration.
+Repeating this over time for a fixed number of restarts with updated parameters required by the inner-loop, defines our outer loop procedure.
+The mechanics of this procedure gives rise to a diagonal equilibrium tracking method (DANTE), formally described in Algorithm 2.
 Algorithm 2 DANTE (DiAgoNal equilibrium Tracking mEthod)
-$w_{0}\in\mathcal{H}$ and $(\tau_{0,k}),(\theta_{0,k})_{k}$ given.
+$w_{0}\in\mathcal{H}$ given initial point. $(\tau_{0,k}),(\theta_{0,k})_{k}$ given parameters and $N\geq 1$ given number of restarts.
 Set $\lambda_{0}=1$ and $S_{0}=0$.
@@
 for $n=0,\ldots,N-1$ do
-   Set $v=\mathtt{KM}(w_{n},\mathsf{T}^{(w_{n},\beta_{n})}_{n,\bullet},(\tau_{n,k})_{k},(\theta_{n,k})_{k},\varepsilon_{n})$.
+   Set $v=\mathtt{KM}(w_{n},(\mathsf{T}^{(w_{n},\beta_{n})}_{n,k})_{k},(\tau_{n,k})_{k},(\theta_{n,k})_{k},\varepsilon_{n})$.
    Let $w_{n+1}=\mathsf{Z}^{(w_{n},\beta_{n})}(v)$ denote the next anchor point.
@@
 return $\overline{w}_{N}$
-Note that DANTE is a method rather than an algorithm as we do not specify how the parameters required by the fixed-point iteration procedure $\mathtt{KM}(v,\mathsf{T}^{(w,\beta)}_{\bullet},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$ are set.
+Note that DANTE is a method rather than a bona-fide algorithm as we do not specify how the parameters required by the fixed-point iteration $\mathtt{KM}(v,(\mathsf{T}^{(w,\beta)}_{k})_{k},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$ are updated.
 This is a part of the user-defined input and should be decided based on the eventually known problem structure.
@@
 [Cortild:2024aa, Hypothesis 2.1] Assume $(\theta_{k})_{k}\subset(0,\overline{\theta}]$ and that $(\tau_{k})_{k}\subset[0,\overline{\tau}]$ is monotonically non-decreasing, where $\overline{\theta},\overline{\tau}\in(0,1)$.
-Moreover, assume that, for all $k\in\mathbb{N}$,
+Moreover,
 $Q_{k}\tau_{k}(1+\tau_{k})+(\theta_{k}^{-1}-1)\tau_{k}(1-\tau_{k})-Q_{k}(\theta_{k-1}^{-1}-1)(1-\tau_{k-1})<0,$ (3.1)
-where $Q_{k}=1-\theta_{k}(1-q_{k}^{2})$, and that $\overline{Q}\triangleq\sup_{k\in\mathbb{N}}Q_{k}<1$.
-Under the above assumption, we can guarantee strong convergence of the iterates.
+for all $k\in\mathbb{N}$, where $Q_{k}=1-\theta_{k}(1-q_{k}^{2})$, and that $\overline{Q}\triangleq\sup_{k\in\mathbb{N}}Q_{k}<1$.
+Under the above assumption, we can guarantee strong convergence of the iterates at a linear rate.
 Theorem 3.1.
@@
 $\leq 2(1+\overline{\tau})\max(\|v_{k+1}-p\|,\|v_{k}-p\|,\|v^{k-1}-p\|).$
-The result follows by the square-summability of $(v_{k}-p)$, and the rate provided in Theorem 3.1. $\blacksquare$
+The result follows by the square-summability of $(v_{k}-p)_{k}$, and the rate provided in Theorem 3.1. $\blacksquare$
 The rate in Corollary 3.2 does not provide a user-friendly quantitative result unless one can control the quantity $\|v_{0}-p\|$.
@@
 Assumption 4.
-We assume that $\operatorname{dom}(\mathsf{T}_{k})\subset\operatorname{dom}(\mathsf{M})$ for all $k\in\mathbb{N}$.
+We assume that $\operatorname{dom}(\mathsf{T}^{(w,\beta)}_{k})\subset\operatorname{dom}(\mathsf{M})$ for all $k\in\mathbb{N}$ and all $(w,\beta)$.
 Remark 3.1.
 Assumption 4 might seem restrictive, but in fact, they are satisfied by many operators.
-For instance, if $\mathsf{M}=\mathsf{F}+\operatorname{\mathsf{NC}}_{\mathcal{K}}$, where $\operatorname{\mathsf{NC}}_{\mathcal{K}}$ is the normal cone of ${\mathcal{K}}=\operatorname{dom}(\mathsf{M})$ and $\mathsf{F}$ is a Lipschitz monotone operator, the forward-backward operator (FB) satisfies the desired conditions, as the resolvent of $\operatorname{\mathsf{NC}}_{\mathcal{K}}$ reduces to a projection onto $\mathcal{K}$}, such that $\mathsf{T}(x)\in{\mathcal{K}}$ for all $x\in\mathcal{H}$.
-The forward-backward operator and other examples are discussed in Section 4.
+As a concrete example, assume that we know that the lower level is formulated in terms of the operator $\mathsf{M}=\mathsf{F}+\operatorname{\mathsf{NC}}_{\mathcal{K}}$, where $\operatorname{\mathsf{NC}}_{\mathcal{K}}$ is the normal cone of the closed convex set ${\mathcal{K}}=\operatorname{dom}(\mathsf{M})$ and $\mathsf{F}:\mathcal{H}\to\mathcal{H}$ is a Lipschitz continuous monotone operator.
+With this problem structure, a powerful fixed-point encoding strategy is the forward-backward operator.
+Adapted to the auxiliary problem (Aux), this method is defined in terms of the mapping
+$\mathsf{T}^{(w,\beta)}_{k}(x)=\Pi_{\mathcal{K}}(x-\gamma_{k}(\mathsf{F}(x)+\beta\mathsf{G}(x)+\alpha(x-w))),$
+$\gamma_{k}>0$ being a suitably chosen step-size parameter.
+Clearly, $\mathsf{T}^{(w,\beta)}_{k}(x)\in{\mathcal{K}}$ for all $x\in\mathcal{H}$.
+The forward-backward operator and other examples are discussed in detail in Section 4.
 Under these assumptions, we can achieve a user-friendly rate, as presented in the following corollary.
 Corollary $3.$
-Assume Assumptions 2, 3, and 4 hold. Then
+Let Assumptions 2-4 be in place. Then,
 $\|v_{k+1}-z_{k}\|\leq C\cdot\overline{Q}^{k/2},$
 for all $k\geq 1$, where
-$C\triangleq\frac{2(1+\overline{\tau})}{\sqrt{\overline{Q}(1-\overline{\tau})(1-\overline{\theta})}}\cdot\operatorname{diam}(\operatorname{dom}(\mathsf{M})).$ (Def-$C$)
+$C\triangleq\frac{2(1+\overline{\tau})}{\sqrt{\overline{Q}(1-\overline{\tau})(1-\overline{\theta})}}\cdot\operatorname{diam}(\operatorname{dom}(\mathsf{M})).$ (3.2)
 Specifically,
-$\mathbb{K}(\varepsilon)\leq\overline{K}_{\varepsilon}\triangleq\left\lceil 2\log(C/\varepsilon)/\log(1/\overline{Q})\right\rceil.$
+$\mathbb{K}(\varepsilon)\leq\overline{K}_{\varepsilon}\triangleq\left\lceil 2\log(C/\varepsilon)/\log(1/\overline{Q})\right\rceil.$ (3.3)
 Proof.
@@
 Remark 3.2.
-By [Cortild:2024aa, Theorem 2.1], we also know that weak convergence of the iterates $(v_{k})$ and the velocities $z_{k}$ is guaranteed when $\mathsf{T}_{k}$ is just nonexpansive.
-Moreover, if $\mathsf{T}_{k}$ is a fixed-point encoding of Problem (Aux), which has a unique solution $p$ by construction, we know that the sequence convergences weakly to this unique fixed-point $p$.
+By [Cortild:2024aa, Theorem 2.1], we also know that weak convergence of the iterates $(v_{k})$ and the velocities $(z_{k})$ is guaranteed when $\mathsf{T}_{k}$ is just nonexpansive.
+Moreover, if $\mathsf{T}_{k}$ is a fixed-point encoding of Problem (Aux), which has a unique solution $p$ by construction, we know that the sequence convergences weakly to this unique fixed-point.
 Specifically, in the finite-dimensional setting, we know that $\|v_{k}-p\|\to 0$ and $\|v_{k+1}-z_{k}\|\to 0$, meaning that Algorithm 2 still can be applied.
-It is however important to note that Lemma 2.1 no longer holds, and that we do not have an appropriate stopping criterion to guarantee $\|w^{+}-\overline{u}_{\beta}(w)\|\leq\mathtt{e}$.
-Within our final numerical implementation in Section 5.3, we allow for nonexpansive operators, corresponding to $q=1$, heavily noting that the theoretical guarantees are weakened in this setting.
+However, Lemma 2.1 no longer holds, and we do not have an appropriate stopping criterion to guarantee $\|w^{+}-\overline{u}_{\beta}(w)\|\leq\mathtt{e}$.
+Within our final numerical implementation in Section 5.3, we experimentally test our algorithm on problems displaying only nonexpansive operators, corresponding to $q=1$, noting that the theoretical guarantees are weakened in this setting.
 Remark 3.3.
@@
 As such, Corollary 3.2 still holds, but again without rate guarantees.
-Specifically, we have no apriori upper bound on the required number of iterations required for Algorithm 1 to terminate, but we know it will terminate.
+Specifically, we have no a-priori upper bound on the required number of iterations required for Algorithm 1 to terminate, but we know it will terminate.
 3.1.1 Parameters Analysis
@@
 3.2 Convergence Analysis of Restarting Procedure
-In the following, we assume that the operator $\mathsf{G}$ is $\mu$-strongly monotone with $\mu\geq 0$.
-The case in which $\mathsf{G}$ is merely monotone is recovered by setting $\mu=0$ throughout the subsequent analysis.
-We introduce two gap functions, one associated with optimality and one with feasibility.
+For our complexity analysis, we introduce two gap functions, one associated with the solution set $\mathcal{S}_{0}$, and one with feasibility $\operatorname{Zer}(\mathsf{M})$.
+Definition 3.4.
 The optimality gap is defined as
-$\operatorname{Gap}_{\operatorname{opt}}(u)\triangleq\sup_{v\in\mathcal{S}_{0}}\langle\mathsf{G}v,u-v\rangle.$ ($\operatorname{Gap}_{\operatorname{opt}}$)
+$\operatorname{Gap}_{\operatorname{opt}}(u)\triangleq\sup_{v\in\mathcal{S}_{0}}\langle\mathsf{G}v,u-v\rangle.$ (3.4)
 We observe that if $u$ is feasible, namely if $u\in\mathcal{S}_{0}=\operatorname{Zer}(\mathsf{M})$, then the condition $\operatorname{Gap}_{\operatorname{opt}}(u)\leq 0$ is equivalent to $u\in\operatorname{Zer}(\mathsf{G}+\operatorname{\mathsf{NC}}_{\mathcal{S}_{0}})$, such that $u$ solves Problem (P).
@@
 $x\in\mathcal{S}_{0}=\operatorname{Zer}(\mathsf{M})\Leftrightarrow\inf_{y\in\operatorname{dom}(\mathsf{M})}\inf_{v\in\mathsf{M}y}\langle 0-v,x-y\rangle\geq 0.$
-As such, we define the feasibility gap $\operatorname{Gap}_{\operatorname{feas}}:\mathcal{H}\to\mathbb{R}$ as
-$\operatorname{Gap}_{\operatorname{feas}}(x)=\sup_{y\in\operatorname{dom}(\mathsf{M}),v\in\mathsf{M}y}\langle v,x-y\rangle.$ ($\operatorname{Gap}_{\operatorname{feas}}$)
+Definition 3.5.
+The feasibility gap $\operatorname{Gap}_{\operatorname{feas}}:\mathcal{H}\to\mathbb{R}$ is defined as
+$\operatorname{Gap}_{\operatorname{feas}}(x)=\sup_{y\in\operatorname{dom}(\mathsf{M}),v\in\mathsf{M}y}\langle v,x-y\rangle.$ (3.5)
 We note that $\operatorname{Gap}_{\operatorname{feas}}(x)\leq 0$ implies $x\in\operatorname{Zer}(\mathsf{M})$, and that $\operatorname{Gap}_{\operatorname{feas}}(x)\geq 0$ holds for all $x\in\mathcal{H}$ since, by assumption, the problem admits at least one solution; in particular, $\operatorname{Zer}(\mathsf{M})\neq\varnothing$.
-Since $\operatorname{Gap}_{\operatorname{opt}}$ and $\operatorname{Gap}_{\operatorname{feas}}$ are pointwise suprema of affine functions, both are weakly lower semicontinuous.
+Since $\operatorname{Gap}_{\operatorname{opt}}$ and $\operatorname{Gap}_{\operatorname{feas}}$ are pointwise suprema of affine functions, both are weakly lower semicontinuous and convex.
 Our convergence rates in terms of the optimality gap require a lower bound on the a priori signless optimality gap function $\operatorname{Gap}_{\operatorname{opt}}$.
@@
 An important class of examples arises when $\mathsf{M}$ is single-valued and $\rho=1$, notably in monotone linear complementarity problems in finite dimensions under nondegeneracy conditions [burke_weak_1993, pang1997error].
+Remark 3.4.
 In the case where the lower-level problem is determined by a convex optimization problem, weak-sharpness implies an Hölderian errror bound, an assumption already imposed by [Cabot2005] in the context of hierarchical minimization.
-Specifically, let us assume that $\mathsf{M}=\partial f$ for a proper convex and lower semi-continuous function $f:\mathcal{H}\to\mathbb{R}\cup\{+\infty\}$.
+Specifically, let us assume that $\mathsf{M}=\partial f$ the subgradient of a proper convex and lower semi-continuous function $f:\mathcal{H}\to\mathbb{R}\cup\{+\infty\}$.
 Then,
@@
 We next define the following quantities, which are finite because $\operatorname{dom}(\mathsf{M})$ is assumed to be bounded:
-$D_{\mathsf{M}}=\operatorname{diam}(\operatorname{dom}(\mathsf{M})),\hskip 17.00024pt\hskip 17.00024pt$ $C_{\mathsf{M}}=\sup_{x\in\operatorname{dom}(\mathsf{M})}\sup_{v\in\mathsf{M}x}\|v\|,\qquad$ $C_{\mathsf{G}}=\sup_{x\in\operatorname{dom}(\mathsf{M})}\mathsf{G}x.$
+$D_{\mathsf{M}}=\operatorname{diam}(\operatorname{dom}(\mathsf{M})),\hskip 17.00024pt\hskip 17.00024pt$ $C_{\mathsf{M}}=\sup_{x\in\operatorname{dom}(\mathsf{M})}\sup_{v\in\mathsf{M}x}\|v\|,\qquad$ $C_{\mathsf{G}}=\sup_{x\in\operatorname{dom}(\mathsf{M})}\mathsf{G}(x).$
 We also give more explicit expressions for the averaging coefficients $(\lambda_{n})_{n}$ and the corresponding averaged iterates which are produced by DANTE. They read as
@@
 Assumption 5.
-We assume the following on the parameter sequences:
+The parameter sequences satisfy the following conditions:
   1. 1.
@@
 Proposition 3.7.
-For any $x\in\operatorname{dom}(\mathsf{M})$ and $v\in\mathsf{M}x$, it holds that
-$\lambda_{n}\langle v,w_{n+1}-x\rangle+\lambda_{n}\beta_{n}\langle\mathsf{G}x,w_{n+1}-x\rangle\leq-\frac{\alpha\lambda_{n+1}}{2}\|w_{n+1}-x\|^{2}+\frac{\alpha\lambda_{n}}{2}\|w_{n}-x\|^{2}+C_{1}\lambda_{n}\mathtt{e}_{n},$ (3.4)
+For any $(x,v)\in\operatorname{Gr}(\mathsf{M})$, it holds that
+$\lambda_{n}\langle v,w_{n+1}-x\rangle+\lambda_{n}\beta_{n}\langle\mathsf{G}(x),w_{n+1}-x\rangle\leq-\frac{\alpha\lambda_{n+1}}{2}\|w_{n+1}-x\|^{2}+\frac{\alpha\lambda_{n}}{2}\|w_{n}-x\|^{2}+C_{1}\lambda_{n}\mathtt{e}_{n},$ (3.8)
 where
@@
 $\frac{1}{2}\|w_{n+1}-x\|^{2}-\frac{1}{2}\|w_{n}-x\|^{2}\leq\langle w_{n}-\overline{u}_{n+1},x-\overline{u}_{n+1}\rangle+2(D_{\mathsf{M}}+\mathtt{e}_{\max})\mathtt{e}_{n}.$
-As $\beta_{n}\mathsf{G}\overline{u}_{n+1}-\alpha(\overline{u}_{n+1}-w_{n})\in\mathsf{M}\overline{u}_{n+1}$ and $v\in\mathsf{M}x$, it follows by monotonicity of $\mathsf{M}$ and by $\mu$-monotonicity of $\mathsf{G}$ that
-$\langle w_{n}-\overline{u}_{n+1},x-\overline{u}_{n+1}\rangle$ $\leq\frac{1}{\alpha}\langle v,x-\overline{u}_{n+1}\rangle+\frac{\beta_{n}}{\alpha}\langle\mathsf{G}\overline{u}_{n+1},x-\overline{u}_{n+1}\rangle$
-$\leq\frac{1}{\alpha}\langle v,x-\overline{u}_{n+1}\rangle+\frac{\beta_{n}}{\alpha}\langle\mathsf{G}x,x-\overline{u}_{n+1}\rangle-\frac{\mu\beta_{n}}{\alpha}\|x-\overline{u}_{n+1}\|^{2}.$
+As $\beta_{n}\mathsf{G}(\overline{u}_{n+1})-\alpha(\overline{u}_{n+1}-w_{n})\in\mathsf{M}(\overline{u}_{n+1})$ and $v\in\mathsf{M}(x)$, it follows by monotonicity of $\mathsf{M}$ and by $\mu$-monotonicity of $\mathsf{G}$ that
+$\langle w_{n}-\overline{u}_{n+1},x-\overline{u}_{n+1}\rangle$ $\leq\frac{1}{\alpha}\langle v,x-\overline{u}_{n+1}\rangle+\frac{\beta_{n}}{\alpha}\langle\mathsf{G}(\overline{u}_{n+1}),x-\overline{u}_{n+1}\rangle$
+$\leq\frac{1}{\alpha}\langle v,x-\overline{u}_{n+1}\rangle+\frac{\beta_{n}}{\alpha}\langle\mathsf{G}(x),x-\overline{u}_{n+1}\rangle-\frac{\mu\beta_{n}}{\alpha}\|x-\overline{u}_{n+1}\|^{2}.$
 Now we note that
@@
 Combining the previous three relationships yields
-$\frac{1}{2}\|w_{n+1}-x\|^{2}-\frac{1}{2}\|w_{n}-x\|^{2}$ $\leq\frac{1}{\alpha}\langle v,x-\overline{u}_{n+1}\rangle+\frac{\beta_{n}}{\alpha}\langle\mathsf{G}x,x-\overline{u}_{n+1}\rangle-\frac{\mu\beta_{n}}{\alpha}\|{x-w_{n+1}}\|^{2}$
+$\frac{1}{2}\|w_{n+1}-x\|^{2}-\frac{1}{2}\|w_{n}-x\|^{2}$ $\leq\frac{1}{\alpha}\langle v,x-\overline{u}_{n+1}\rangle+\frac{\beta_{n}}{\alpha}\langle\mathsf{G}(x),x-\overline{u}_{n+1}\rangle-\frac{\mu\beta_{n}}{\alpha}\|{x-w_{n+1}}\|^{2}$
 $\qquad+2\frac{\mu\beta_{n}}{\alpha}(D_{\mathsf{M}}+\mathtt{e}_{max})\mathtt{e}_{n}+2(D_{\mathsf{M}}+\mathtt{e}_{\max})\mathtt{e}_{n}.$
 Rearranging the terms and bounding $\beta_{n}\leq\beta_{\max}$ gives
-$\left(\frac{1}{2}+\frac{\mu\beta_{n}}{\alpha}\right)\|w_{n+1}-x\|^{2}-\frac{1}{2}\|w_{n}-x\|^{2}\leq\frac{1}{\alpha}\langle v,x-\overline{u}_{n+1}\rangle+\frac{\beta_{n}}{\alpha}\langle\mathsf{G}x,x-\overline{u}_{n+1}\rangle+2\left(1+\frac{\mu\beta_{\max}}{\alpha}\right)(D_{\mathsf{M}}+\mathtt{e}_{\max})\mathtt{e}_{n}.$
+$\left(\frac{1}{2}+\frac{\mu\beta_{n}}{\alpha}\right)\|w_{n+1}-x\|^{2}-\frac{1}{2}\|w_{n}-x\|^{2}\leq\frac{1}{\alpha}\langle v,x-\overline{u}_{n+1}\rangle+\frac{\beta_{n}}{\alpha}\langle\mathsf{G}(x),x-\overline{u}_{n+1}\rangle+2\left(1+\frac{\mu\beta_{\max}}{\alpha}\right)(D_{\mathsf{M}}+\mathtt{e}_{\max})\mathtt{e}_{n}.$
 After noting that
@@
 and
-$\langle\mathsf{G}x,x-\overline{u}_{n+1}\rangle=\langle\mathsf{G}x,x-w_{n+1}\rangle+\langle\mathsf{G}x,w_{n+1}-\overline{u}_{n+1}\rangle\leq\langle\mathsf{G}x,x-w_{n+1}\rangle+C_{\mathsf{G}}\mathtt{e}_{n},$
+$\langle\mathsf{G}(x),x-\overline{u}_{n+1}\rangle=\langle\mathsf{G}(x),x-w_{n+1}\rangle+\langle\mathsf{G}(x),w_{n+1}-\overline{u}_{n+1}\rangle\leq\langle\mathsf{G}(x),x-w_{n+1}\rangle+C_{\mathsf{G}}\mathtt{e}_{n},$
 and multiplying both sides by $\lambda_{n}\alpha$, we obtain
-$\lambda_{n}\langle v,w_{n+1}-x\rangle$ $+\lambda_{n}\beta_{n}\langle\mathsf{G}x,w_{n+1}-x\rangle$
+$\lambda_{n}\langle v,w_{n+1}-x\rangle$ $+\lambda_{n}\beta_{n}\langle\mathsf{G}(x),w_{n+1}-x\rangle$
 $\leq-\left(\frac{1}{2}+\frac{\mu\beta_{n}}{\alpha}\right)\lambda_{n}\alpha\|w_{n+1}-x\|^{2}+\frac{\lambda_{n}\alpha}{2}\|{w_{n}-x}\|^{2}$
@@
 Select $x\in\operatorname{Zer}(\mathsf{M})$ and $v=0$ in (3.8), and sum the result for $n=0,\ldots,N-1$ to obtain that
-$\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}\langle\mathsf{G}x,w_{n+1}-x\rangle$ $\leq\frac{-\alpha\lambda_{N}}{2}\|w_{N}-x\|^{2}+\frac{\alpha\lambda_{0}}{2}\|w_{0}-x\|^{2}+C_{1}\cdot\sum_{n=0}^{N-1}\lambda_{n}\mathtt{e}_{n}$
+$\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}\langle\mathsf{G}(x),w_{n+1}-x\rangle$ $\leq\frac{-\alpha\lambda_{N}}{2}\|w_{N}-x\|^{2}+\frac{\alpha\lambda_{0}}{2}\|w_{0}-x\|^{2}+C_{1}\cdot\sum_{n=0}^{N-1}\lambda_{n}\mathtt{e}_{n}$
 $\leq{\alpha\lambda_{0}}(\|w_{0}-\overline{u}_{0}\|^{2}+\|\overline{u}_{0}-x\|^{2})+C_{1}\cdot\sum_{n=0}^{N-1}\lambda_{n}\mathtt{e}_{n}$
@@
 After normalization, this yields
-$\langle\mathsf{G}x,\overline{w}_{N}-x\rangle\leq C_{2}\cdot\frac{1}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}+C_{1}\cdot\frac{\sum_{n=0}^{N-1}\lambda_{n}\mathtt{e}_{n}}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}.$
+$\langle\mathsf{G}(x),\overline{w}_{N}-x\rangle\leq C_{2}\cdot\frac{1}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}+C_{1}\cdot\frac{\sum_{n=0}^{N-1}\lambda_{n}\mathtt{e}_{n}}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}.$
 The inequality follows by taking the supremum over $x\in\operatorname{Zer}(\mathsf{M})$.
@@
 Multiply both sides of (3.8) by $\beta_{n}$, and recall that $\beta_{n+1}\leq\beta_{n}$. Moreover, notice that
-$\langle\mathsf{G}x,w_{n+1}-x\rangle\geq-\|{\mathsf{G}x}\|\cdot\|{w_{n+1}-x}\|\geq-\|\mathsf{G}x\|\cdot(\|w_{n+1}-\overline{u}_{n+1}\|+\|\overline{u}_{n+1}-x\|)\geq-C_{\mathsf{G}}(\mathtt{e}_{n}+D_{\mathsf{M}}).$
-It then follows $\langle\mathsf{G}x,w_{n+1}-x\rangle\leq C_{\mathsf{G}}(\mathtt{e}_{\max}+D_{\mathsf{M}})$. Summing for $n=0,\ldots,N-1$ then yields
+$\langle\mathsf{G}(x),w_{n+1}-x\rangle\geq-\|{\mathsf{G}(x)}\|\cdot\|{w_{n+1}-x}\|\geq-\|\mathsf{G}(x)\|\cdot(\|w_{n+1}-\overline{u}_{n+1}\|+\|\overline{u}_{n+1}-x\|)\geq-C_{\mathsf{G}}(\mathtt{e}_{n}+D_{\mathsf{M}}).$
+It then follows $\langle\mathsf{G}(x),w_{n+1}-x\rangle\leq C_{\mathsf{G}}(\mathtt{e}_{\max}+D_{\mathsf{M}})$. Summing for $n=0,\ldots,N-1$ then yields
 $\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}\langle v,w_{n+1}-x\rangle$ $\leq-\frac{\alpha\lambda_{N}\beta_{N}}{2}\|w_{N}-x\|^{2}+\frac{\alpha\lambda_{0}\beta_{0}}{2}\|w_{0}-x\|^{2}+C_{1}\cdot\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}\mathtt{e}_{n}$
@@
 $\blacksquare$
-We may now combine the two previous lemmas to state our main theorem.
+We may now combine the two previous lemmas to state our main theorem of this paper.
+We establish a combined complexity result in terms of the two introduced gap functions using the averaged trajectory $\overline{w}_{N}$.
 Theorem 3.10 (Main Theorem).
@@
     The upper bound is just Lemma 3.8.
-For the lower bound, we invoke the Cauchy-Schwarz inequality $\langle\mathsf{G}x,\overline{w}_{N}-x\rangle\geq\|{\mathsf{G}x}\|\cdot\|{\overline{w}_{N}-x}\|$, implying
+For the lower bound, we invoke the Cauchy-Schwarz inequality $\langle\mathsf{G}(x),\overline{w}_{N}-x\rangle\geq\|{\mathsf{G}(x)}\|\cdot\|{\overline{w}_{N}-x}\|$, implying
     $\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{N})\geq-C_{\mathsf{G}}\operatorname{dist}(\overline{w}_{N},\mathcal{S}_{0}).$
@@
 Corollary 3.11.
-Let $(w_{n})$ be generated through Algorithm 2 when $\mathsf{G}$ is merely monotone (i.e.
+Let $(w_{n})_{n=0}^{N}$ be generated through Algorithm 2.
+Assume that $\mathsf{G}$ is merely monotone (i.e.
 $\mu=0$).
@@
 $0\leq\operatorname{Gap}_{\operatorname{feas}}(\overline{w}_{N})\leq\mathcal{O}\left(\frac{\log(N)}{N}\right).$
-If, moreover, $\operatorname{Zer}(\mathsf{M})$ is weakly sharp with constant $\kappa$ of order $\rho\geq 1$, then it also holds that $\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{N})\geq\mathcal{O}(\operatorname{Gap}_{\operatorname{feas}}(\overline{w}_{N})^{1/\rho}).$
+If, moreover, $\mathcal{S}_{0}$ is weakly sharp with constant $\kappa$ of order $\rho\geq 1$, then it also holds that $\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{N})\geq\mathcal{O}(\operatorname{Gap}_{\operatorname{feas}}(\overline{w}_{N})^{1/\rho}).$
 Proof.
@@
 The conclusion follows by applying these estimates to Theorem 3.10. $\blacksquare$
-3.3 Combined Complexity Analysis
-We now finally perform the complete complexity analysis of our scheme DANTE by combining our bounds derived for Algorithms 1 and 2.
-Before stating the final convergence result, we make one final assumption.
+3.3 Total Complexity Analysis
+We now finally perform the complete complexity analysis of our scheme DANTE by combining the bounds derived for Algorithms 1 and 2.
+Given a sequences of user-defined tolerances $(\varepsilon_{n})_{n=0}^{N}$, we known that Algorithm 1 terminates after at most $\overline{K}_{\varepsilon_{n}}\equiv\overline{K}_{n}$ iterations (Corollary 3.3, eq.
+(3.3)).
+To assess the total iteration complexity of Algorithm 2 for a pre-defined number of restarts $N\geq 1$, we thus need to compute the total iteration complexity
+$\mathcal{C}(N,(\varepsilon_{n})_{n=0}^{N})\triangleq\sum_{n=0}^{N-1}\mathbb{K}(\varepsilon_{n}).$ (3.14)
+To estimate the total iteration complexity, we make one final assumption.
 Assumption 6.
-We assume $\mathsf{Z}_{n}$ is domain-forward on $\operatorname{dom}(\mathsf{M})$, namely that $\mathsf{Z}_{n}(\operatorname{dom}(\mathsf{M}))\subset\operatorname{dom}(\mathsf{M})$.
+We assume $\mathsf{Z}_{n}$ is domain-forward on $\operatorname{dom}(\mathsf{M})$:
+$\mathsf{Z}_{n}(\operatorname{dom}(\mathsf{M}))\subset\operatorname{dom}(\mathsf{M}).$
 Finally, we assume that $\theta_{n,k}$ is uniformly lower bounded by a positive quantity, such that $\overline{Q}_{n}\leq\overline{Q}<1$.
@@
 Theorem 3.13.
-Assume the setting of Corollary 3.3 and that Assumptions 2, 3, 4 and 6 hold. Algorithm 2 (DANTE) reaches a point $\overline{w}_{N}$ satisfying the inequalities in Theorem 3.8 after at most
-$K=\frac{N(2\log(C)+\log(\overline{Q}))+2\sum_{n=1}^{N}\log(1/\varepsilon_{n})}{\log(1/\overline{Q})}$
-inner iterations. If $\varepsilon_{n}=(n+1)^{-\eta}$ where $\eta\geq b$, we reach the desired point $\overline{w}_{N}$ after at most ${\mathcal{O}}(N\log N)$ inner iterations.
+Let Assumptions 1-6 hold. Given $N\geq 1$, and inner loop tolerance sequence $(\varepsilon_{n})_{n=0}^{N}$, we have
+$\mathcal{C}(N,(\varepsilon_{n})_{n=0}^{N})\leq\frac{N(2\log(C)+\log(\overline{Q}))+2\sum_{n=1}^{N}\log(1/\varepsilon_{n})}{\log(1/\overline{Q})},$
+where the constant $C$ is defined in (3.2) and $\overline{Q}$ is defined in Assumption 3.
+Specifically, if $\varepsilon_{n}=(n+1)^{-\eta}$ and $\beta_{n}=(n+1)^{-b}$ where $\eta\geq b$, then $\mathcal{C}(N,(\varepsilon_{n})_{n=0}^{N})={\mathcal{O}}(N\log N)$.
 Proof.
-By Corollary 3.3, each outer loop $n\in\{0,\ldots,N-1\}$ runs at most
-$\overline{K}_{n}=\left\lceil\frac{2\log(C)+2\log(1/\varepsilon_{n})}{\log(1/\overline{Q})}\right\rceil\leq\frac{2\log(C)+2\log(1/\varepsilon_{n})}{\log(1/\overline{Q})}+1$
-inner iterations. Specifically, the inequalities of Theorem 3.8 are satisfied after at most
-$K=\sum_{n=0}^{N-1}\overline{K}_{n}\leq\frac{2N\log(C)+2\sum_{n=1}^{N}\log(1/\varepsilon_{n})}{\log(1/\overline{Q})}+N=\mathcal{O}\left(N+\sum_{n=0}^{N-1}\log(1/\varepsilon_{n})\right)$
-By setting $\varepsilon_{n}=(n+1)^{-\eta}$, we obtain that $\sum_{n=0}^{N-1}\log(1/\varepsilon_{n})\leq\eta N\log(N)$, as wanted. $\blacksquare$
-Remark 3.5.
+By Corollary 3.3, for each restarting epoch $n\in\{0,\ldots,N-1\}$ the total number of iterations of the inner loop is bounded by
+$\mathbb{K}(\varepsilon_{n})\leq\overline{K}_{n}=\left\lceil\frac{2\log(C)+2\log(1/\varepsilon_{n})}{\log(1/\overline{Q})}\right\rceil\leq\frac{2\log(C)+2\log(1/\varepsilon_{n})}{\log(1/\overline{Q})}+1.$
+Hence,
+$\sum_{n=0}^{N-1}\overline{K}_{n}\leq\frac{2N\log(C)+2\sum_{n=1}^{N}\log(1/\varepsilon_{n})}{\log(1/\overline{Q})}+N=\mathcal{O}\left(N+\sum_{n=0}^{N-1}\log(1/\varepsilon_{n})\right)$
+gives an upper bound on the total number of iterations for Algorithm 2 for $N\geq 1$ restarts.
+By setting $\varepsilon_{n}=(n+1)^{-\eta}$, we obtain that $\sum_{n=0}^{N-1}\log(1/\varepsilon_{n})\leq\eta N\log(N)$.
+$\blacksquare$
 Theorem 3.13 shows that accounting for the total number of inner iterations, rather than only the number of outer iterations, introduces merely a logarithmic factor.
 This is expected, as double-loop schemes inherently incur such a term.
-Consequently, the complexities stated in Corollaries 3.9 and 3.10 are preserved, up to logarithmic factors, when measured in terms of total inner iterations.
+Consequently, the complexities stated in Corollaries 3.11 and 3.12 for the gap functions remain intact, up to logarithmic factors, when accounting for the total number of inner iterations.
 4 Relevant Splitting Methods
@@
 Remark 4.1.
-With the following examples it will become clear that the fixed-point encoding $\mathsf{T}^{(w,\beta)}_{k}$ not only depends on the problem parameters $\alpha$ and $\beta$ but also on the step size $\gamma_{k}$.
-Therefore, we treat the step size policy is treated in our approach as an ingredient of the fixed-point encoding strategy and consequently incorporated in the map $\mathsf{T}^{(w,\beta)}_{k}$, rather than being a parameter of the inner loop presented in Algorithm 1.
+With the following examples, it will become clear that the fixed-point encoding $\mathsf{T}^{(w,\beta)}_{k}$ not only depends on the problem parameters $\alpha$ and $\beta$ but also on step size sequences $(\gamma_{k})_{k}$.
+Therefore, we treat the step size policy as an ingredient of the fixed-point encoding strategy and consequently incorporate it in the map $\mathsf{T}^{(w,\beta)}_{k}$, rather than treating it as a parameter of the inner loop presented in Algorithm 1.
 4.1 Two-Operator Splitting Schemes
@@
   1. 1.
-    The lower level solution set $\mathcal{S}_{0}$ is nonempty;
+    The lower-level solution set $\mathcal{S}_{0}$ is nonempty;
   2. 2.
@@
 $\mathsf{T}^{(w,\beta)}_{\gamma}=\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}\circ(\operatorname{Id}-\gamma\Phi_{\alpha,\beta}^{w}).$ (FB)
-We note that $\operatorname{Fix}(\mathsf{T}^{(w,\beta)}_{\gamma})=\operatorname{Zer}(A+\Phi_{\alpha,\beta}^{w})$, so $\mathsf{Z}=\operatorname{Id}$ is the fixed-point transportation map.
+We note that $\operatorname{Fix}(\mathsf{T}^{(w,\beta)}_{\gamma})=\operatorname{Zer}(\mathsf{A}+\Phi_{\alpha,\beta}^{w})$, so $\mathsf{Z}=\operatorname{Id}$ is the fixed-point transportation map.
 This is the setting studied in [marschner_tikhonov_2025].
@@
 Proof.
-Pick $z_{1},z_{2}\in\mathcal{H}$.
+Pick $z_{1},z_{2}\in\mathcal{H}$ and set $T=T^{(w,\beta)}_{\gamma}$.
 By nonexpansiveness of $\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}$, and $L_{\alpha,\beta}$-Lipschitz continuity and $\alpha$-strong monotonicity of $\Phi_{\alpha,\beta}$,
@@
 Backward-forward splitting.
-In a similar way, we can employ the Backward-Forward splitting [attouch_backward_2018] for the auxiliary problem. Consider the operator
+Similarly, we can employ the backward-forward splitting [attouch_backward_2018] for the auxiliary problem. Consider the operator
 $\mathsf{T}^{(w,\beta)}_{\gamma}=(\operatorname{Id}-\gamma\Phi_{\alpha,\beta}^{w})\circ\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}.$ (BF)
-Note that $\operatorname{Zer}(\mathsf{A}+\Phi_{\alpha,\beta}^{w})=\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}(\text{Fix}(\mathsf{T}^{(w,\beta)}_{\gamma}))$, which means that the last iterate of the inner loop scheme needs one more evaluation of the resolvent of the operator $\mathsf{A}$ in order to get close to the unique zero of $\mathsf{A}+\Phi_{\alpha,\beta}^{w}$.
+Note that $\operatorname{Zer}(\mathsf{A}+\Phi_{\alpha,\beta}^{w})=\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}(\text{Fix}(\mathsf{T}^{(w,\beta)}_{\gamma}))$, which means that the last iterate of the inner loop scheme needs one more evaluation of the resolvent of the operator $\mathsf{A}$ to get close to the unique zero of $\mathsf{A}+\Phi_{\alpha,\beta}^{w}$.
 Specifically, the fixed-point transportation map is $\mathsf{Z}=\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}$.
 Lemma 4.2.
-Assume Assumption 7 holds and let $\mathsf{T}_{\gamma}$ be the fixed-point encoding defined in (BF). It holds that $\mathsf{T}_{\gamma}$ is $q$-Lipschitz where
+Assume Assumption 7 holds, and let $\mathsf{T}_{\gamma}$ be the fixed-point encoding defined in (BF). It holds that $\mathsf{T}_{\gamma}$ is $q$-Lipschitz where
 $q=\sqrt{1-\gamma(2\alpha-\gamma L^{2}_{\alpha,\beta})},$
@@
 Specifically, we consider an equilibrium problem in Section 5.1, a least-norm least-squares problem in Section 5.2, and an image inpainting problem in Section 5.3.
-All the code may be found on the author’s GitHub page [^1] See https://github.com/Hierarchical-VIs/Regularisation-Methods-for-HVIs., and is run on Intel Xeon Platinum 8380 CPUs.
+All the code may be found on the author’s GitHub page [^1] See https://github.com/Hierarchical-VIs/Regularization-Methods-for-HVIs., and is run on Intel Xeon Platinum 8380 CPUs.
 5.1 Equilibrium selection
@@
 $\langle\nabla\phi(x_{1}^{*},x_{2}^{*}),(w_{1},w_{2})-(x_{1}^{*},x_{2}^{*})\rangle\geq 0\qquad\text{for all $(w_{1},w_{2})\in\operatorname{Zer}(F+\operatorname{\mathsf{NC}}_{X})$}.$
-This matches Problem (P) with $\mathsf{G}=\nabla\phi$ for the upper level operator, and $\mathsf{M}=\mathsf{F}+\operatorname{\mathsf{NC}}_{X}$ for the lower level operator.
+This matches Problem (P) with $\mathsf{G}=\nabla\phi$ for the upper-level operator and $\mathsf{M}=\mathsf{F}+\operatorname{\mathsf{NC}}_{X}$ for the lower-level operator.
 Experimental Setup.
@@
 Moreover, we note that the optimality and feasibility gaps may be computed through a small subproblem in this simple example.
-Figures 3 and 4 plot these values, for the different initial points.
+Figures 3 and 4 plot these values for the different initial points.
 Although we observe differences, we note that all curves are decreasing to 0.
@@
 A vector $s\in\mathbb{R}^{Q}$ is randomly generated with $20$ non-zero entries according to a uniform law on $[0,10]$, and the observation vector $b$ is constructed as $b=As+\nu$, where $\nu$ is a normal random vector with small entries, namely a standard normal vector scaled by a factor $0.1$.
-During the random generation we ensure that $z=A^{\dagger}b\in X$, such that the analytical solution is known to be $z$, where $A^{\dagger}$ is the Moore-Penrose inverse.
+During the random generation, we ensure that $z=A^{\dagger}b\in X$, such that the analytical solution is known to be $z$, where $A^{\dagger}$ is the Moore-Penrose inverse.
 Experimental Setup.
@@
 As expected, we observe a decrease for each problem.
-We note that the initial decrease for the error sequence is consequential and stagnates for a number of iterations afterwards.
+We note that the initial decrease for the error sequence is consequential and stagnates for a number of iterations afterward.
 In fact, for $\alpha=0.1$, the prescribed $2000$ iterations are not enough to observe a second decrease phase, whereas it is sufficient for $\alpha\in\{1,10\}$.
@@
 Experimental Setup.
-For the inner loop, we set $\theta_{k}\equiv\theta=0.75$ and we consider $\tau_{k}\equiv\tau$ to be constantly the largest value satisfying Equation (3.1) across every outer iteration.
+For the inner loop, we set $\theta_{k}\equiv\theta=0.75$ and consider $\tau_{k}\equiv\tau$ to be constantly the largest value satisfying Equation (3.1) across every outer iteration.
 We set $\beta_{n}=(n+1)^{-\eta}$ for $\eta=0.55$ and $\varepsilon_{n}=\overline{\varepsilon}\cdot(n+1)^{-2}$ for $\overline{\varepsilon}=2$.
@@
 Results.
-Figure 9 shows the original image, the corrupt image and the image restored using the above procedure for various proximal parameters.
+Figure 9 shows the original image, the corrupt image, and the image restored using the above procedure for various proximal parameters.
 Table 1 reports the total number of inner iterations run, along with the time spent to produce the results.
@@
 Time in minutes 294 80 38 36
-Figure 10: Evolution of lower-level objective function $\tfrac{1}{2}\|\mathcal{R}(Y)-Y_{\text{corrupt}}\|^{2}+\sigma\|Y\|_{*}$ for various proximal parameters.
+Figure 10: Evolution of lower-level objective function $\tfrac{1}{2}\|\mathcal{R}(w_{n})-Y_{\text{corrupt}}\|^{2}+\sigma\|w_{n}\|_{*}$ for various proximal parameters.
 6 Conclusion
-In this work, we proposed a double loop path following method for solving a hierarchical variational inequalities.
-The inner loop employs a Krasnoselskii-Mann type iteration for solving a suitable auxiliary problem.
+In this work, we propose a double-loop path-following method to solve hierarchical variational inequalities.
+The inner loop aims to solve an auxiliary problem and, given a fixed-point encoding of said problem, employs a Krasnoselskii-Mann-type iteration.
+This step is very flexible regarding the chosen encoding strategy, such that our method includes various known algorithms.
 The outer loop takes the form of a restarting procedure, giving rise to a diagonal equilibrium tracking method (DANTE).
-The obtained convergence rates are comparable with the state of the art results, obtained for specific numerical schemes.
-Our analysis is very flexible with respect to the chosen fixed-point encoding strategy of the lower level solution.
-Importantly, our rates are independent of the specific fixed-point encoding, as long as the stated assumptions are satisfied.
-We belief that our work opens up the possibilities for various interesting directions for future research.
-First, it would be very interesting to device a fully adaptive scheme without restarts.
-Second, an important future extension is the inclusion of stochastic data.
-All these directions are the subject of ongoing research.
-Acknowledgments.
+We obtain weak convergence of the averaged iterates as well as convergence rates, which are comparable to the state-of-the-art results for specific numerical schemes.
+We believe that our work opens up possibilities for various interesting directions for future research.
+First, it would be of practical interest to devise a fully adaptive scheme without restarts.
+Second, an important extension is the inclusion of stochastic data.
+These directions are the subject of ongoing research.
+Acknowledgements
 Daniel Cortild acknowledges the support of the Clarendon Funds Scholarships.